### v1

# Quantum automata and languages of finite index thanks:

Abstract
This paper is a continuation of a previous study conducted in [7, 8] on the so-called measure once finite quantum automata model introduced by Moore and Crutchfield in 2000. We investigate conditions assuring that, given a language recognized by such a device and a language generated by a context-free grammar of finite index or by a matrix context-free grammar, it is recursively decidable whether or not they have a nonempty intersection.
Keywords: Quantum automata, Context-free languages, Algebraic groups, Decidability
AMS subject classifications: 81P68, 68Q45
1 Introduction
In this paper, we investigate an issue of decidability for quantum (finite) automata. Quantum automata were introduced at the beginning of 2000’s in [23] as a new model of language recognizer (cf. [5]). Numerous publications have ever since compared their decision properties to those of other models and, in particular, to probabilistic finite automata. In this theoretical setting, we investigate the decidability of a problem for quantum automata introduced in [6, 7], named Intersection Problem. The starting point of that investigation is a result of Blondel, Jeandel, Koiran, and Portier which shows in [8] that, given a (“measure-once” in the sense of [23]) quantum automaton $\cal Q$, it is decidable whether the language recognized by $\cal Q$ has a non trivial intersection with the free monoid generated by the input alphabet of $\cal Q$ (Emptiness problem).
It is worth remarking that these problems are strictly related to some computational issues concerning matrices. Precisely, a crucial step in the construction of the decision procedure for both the Emptiness and Intersection Problem amounts to effectively compute the Zariski closure of a finitely generated group of matrices over $\mathbb{Q}$. For that task, an appropriate algorithm has been developed in [12] by Derksen, Jeandel, and Koiran.
Before coming back to the Intersection Problem, we find then convenient to recall some developments of the research on that aspect. Two striking results have been obtained in the recent past years in connection with Derksen et al. algorithm. In 2022 Nosan, Pouly, Schmitz, Shirmohammadi, and Worrell, following an alternative approach for such computation, have obtained in [25] a bound on the degree of the polynomials that define the Zariski closure of the set and showed that the latter can be computed in elementary time.
The second was obtained in 2023 by Hrushovski, Ouaknine, Pouly, and Worrel for the study of the effective computation of polynomial invariants for affine programs [17]. Precisely, it is exhibited an algorithm that, given a finite set of square matrices of the same size with coefficients in $\mathbb{Q}$, computes the Zariski closure of the semigroup generated by the set.
It should be pointed out that the decidability issues for matrix semigroups is a challenge problem. Even in small size, a well-known result by Paterson shows that the problem of freeness and that of membership are both undecidable for matrix semigroups over $\mathbb{Z}$ of size $3$ (cf [9]). In contrast, in 2017, a surprising result of Potapov and Semukhin [28] shows that the membership is decidable for non-singular matrix integer semigroups of size $2$. We now come back to the Intersection Problem and give a formal presentation of our work. A finite quantum automaton (in the sense of [23]) is a quadruple
${\cal Q}=(s,\varphi,P,\lambda),$ (1)
where $s\in\mathbb{R}^{n}$ is a row-vector of unit Euclidean norm, $P$ is a projection matrix of $\mathbb{R}^{n}$, $\varphi:\Sigma^{*}\rightarrow O_{n}$ is a morphism from the free monoid $\Sigma^{*}$, generated by the input alphabet $\Sigma$ of $\cal Q$, into the group $O_{n}$ of orthogonal $n\times n$-matrices over $\mathbb{R}$ and the threshold $\lambda$ has value in $\mathbb{R}$.
We are mainly interested in effective properties which require the quantum automaton to be effectively given. We thus consider the model of rational quantum automaton, i.e., where all the coefficients of the components of (1) and $\lambda$ are rational numbers. For a real threshold $\lambda$, the languages recognized by $\cal Q$ with strict and nonstrict threshold $\lambda$ are
$|{\cal Q}_{>}|=\{w\in\Sigma^{*}\mid||s\varphi(w)P||>\lambda\},\quad|{\cal Q}_{\geq}|=\{w\in\Sigma^{*}\mid||s\varphi(w)P||\geq\lambda\},$
where |||||$\cdot$||| | ⋅ | | denotes the Euclidean norm of vectors. Let ${\cal L}$ be a family of effectively defined languages. The problem we tackle is thus the following:
$(L,{\cal Q})$ Intersection
Input: a language $L$ in a family ${\cal L}$ of languages and a rational quantum automaton ${\cal Q}$.
Question: does $L\cap|{\cal Q}_{>}|=\emptyset$ hold?
As already pointed out, it has been proven in [8] the decidability of the problem in the specific case $L$ $=\Sigma^{*}$. This result contrasts with the corresponding problems (strict and non strict, respectively) for probabilistic finite automata. Indeed for this class of automata, the Emptiness problems w.r.t. $|{\cal Q}_{>}|$ and $|{\cal Q}_{\geq}|$ are both undecidable (see [27], Thm 6.17). Also it contrasts with the undecidability of the Emptiness problem of $|{\cal Q}_{\geq}|$ for quantum automata [8]. It is also proven in [21] that both the problems (strict and non strict, respectively) remain undecidable for the “measure many” model of quantum automata, a model not computationally equivalent to the “measure once”, introduced by Kondacs and Watrous in [22].
In order to describe the main contributions of the paper, it may be useful to recall some lines of the technique used in the proof of the Intersection Problem in [7]. For this purpose, following the same strategy of [8], we first observe that the Intersection Problem is equivalent to the inclusion
$\varphi(L)\subseteq|{\cal Q}_{\leq}|.$ (2)
Since the function
$M\rightarrow||sMP||,$
– where $M$ is an arbitrary matrix in $\mathbb{R}^{n\times n}$, and $s,P$ are components of (1) – is continuous, it is sufficient to prove that, for all matrices $M$ in the Euclidean closure ${\mathbf{C}l}(\varphi(L))$ of $\varphi(L)$, the condition $||sMP||\leq\lambda$ holds (see Section 3). Such condition is then effectively tested by running, in parallel, two semialgorithms. The first tests the inclusion $\varphi(L)\subseteq|{\cal Q}_{>}|$ and it is readily seen to be semidecidable. The second, which is the main bulk, is to prove that (2) is semidecidable.
The semidecidability of the latter procedure is achieved by expressing the property (2) in first-order logic of the field of reals, which, in turn, consists to effectively compute a representation of ${\mathbf{C}l}(\varphi(L))$ in term of semialgebraic set (a more general property than algebraic, which allows to cope with the fact that algebraic sets are not closed w.r.t. the operation of product of sets). Afterwards, one applies on the constructed formula the Tarski-Seidenberg quantifier elimination to verify whether (2) holds true or not.
It turns out that the semialgebraicity of ${\mathbf{C}l}(\varphi(L))$, and its effective computation (as a semialgebraic set), are reduced to those of an algebro-combinatorial object that is associated with the grammar $G$ generating $L$, and with every variable $A$ of $G$: the monoid of cycles of $A$. Such monoid, denoted $M_{A}$, corresponds (up to a technical detail) to the matrix image, under taking the morphism $\varphi$ of (1), of the language of cycles in $G$ associated with $A$ (see Eq. (21)).
The reduction process mentioned above is made possible by resorting to two ingredients: the fact that the Euclidean closure ${\mathbf{C}l}(M_{A})$ is an algebraic set (because of a deep result of Algebraic Geometry, namely Theorem 3) and because of a suitably defined effective structuring of the computations, i.e., the derivations, of $G$ (Sections 4.1, 4.2 and 6.1). In this theoretical setting, [7] and this work provides the following picture of the Intersection Problem:
  • If $L$ is a context-free language, then ${\mathbf{C}l}(\varphi(L))$ is semialgebraic. Moreover, if, for every variable $A$ of $G$, ${\mathbf{C}l}(M_{A})$ is effectively computable, then ${\mathbf{C}l}(\varphi(L))$ is so as well (Proposition 9).
  • If, for every variable $A$ of $G$, $M_{A}$ is finitely generated, then ${\mathbf{C}l}(\varphi(L))$ is effectively computable. This is guaranteed by applying Proposition 9 and the results [12, 17, 25] mentioned above. This, in particular, covers the cases of context-free linear languages, bounded semi-linear languages and a new class of languages, proposed in this paper, generated by restricted matrix context-free linear grammars which non trivially extends the previous ones (Proposition 10).
  • The effective computation of ${\mathbf{C}l}(\varphi(L))$ is achieved for a subfamily of finite index context-free languages, called monoidal, by considering the property of (algebraic) irreducibility on the Zariski closure of a family of monoids of cycles canonically associated with $G$ (Proposition 11).
The new results of the paper are the following. In Section 6.1, we extend the main contribution of [7] – i.e., the statement of decidability of the Intersection Problem for linear context-free languages and for bounded semilinear languages – to a class of matrix context-free grammar of finite index. Precisely, whenever the quantum automaton is rational, such result (Proposition 10) allows, on one hand, to recover both the previous two cases as immediate corollaries of such result. On the other, it shows the decidability of the problem for languages of exponential growth that are not context-free. The proof is based upon a combinatorial decomposition of these matrix grammars proposed by Choffrut, Varricchio and the second author in [10] for the study of rational relations. In particular, this decomposition permits a refinement of the algebro-combinatorial machinery developed in [7] for the manipulation of the monoids of cycles of these grammars.
In Section 6.2, we investigate the Intersection Problem for languages defined by finite index context-free grammars. These are grammars $G$ where each word $w$ in the language generated by $G$ is obtained by a derivation $\delta$, whose index is uniformly bounded, i.e., the number of variables in each sentential form of $\delta$ is bounded by an integer not depending on $w$. A remarkable result by Ginsburg and Spanier [14] (see also Nivat [24]) provides a charaterization of such languages in terms of composition of grammars: precisely, each language of this type is generated by a grammar $G$ given by the composition
$G={\cal G}_{1}\circ{\cal G}_{2}\circ\cdots\circ{\cal G}_{k},$ (3)
[…]
We prove that, whatever is the length $k$ of the composition (3) for $G$, if the Zariski closure of the monoids of cycles of the linear grammars of the lowest level ${\cal G}_{k}$ of (3), are (algebraic) irreducible, then one can effectively computes the Euclidean closure ${\mathbf{C}l}(\varphi(L))$ of $\varphi(L)$ (Proposition 11, Corollary 5).
Extensions of such result to broader families of finite index context-free languages seem related to possible extensions of the algorithms [12, 17, 25] (cf Section 7, [$\alpha$]).
The general case of context-free languages is still open and it would be even possible that the problem is undecidable (cf Section 7, [$\beta$]).
The paper is structured in the following way. In Sec. 2 preliminaries on semialgebraic and irreducible algebraic sets are presented. In Section 3 we recall an effective condition to which the decidability of the problem under discussion can be reduced. Section 4 provides combinatorial structurings of context-free languages in the general case and in the finite index case. Section 5 recalls some results relating context-free grammars and semialgebraic sets. In Section 6 the main results of the paper are presented. Section 7 contains some concluding remarks and open problems. Section 8 shows the construction for monoidal grammars on an example.
2 Preliminaries on semialgebraic sets
Let us give first the definition of quantum automaton (cf. [6, 8, 23]).
Definition 1
A quantum automaton ${\cal Q}$ is a quadruple $(s,\varphi,P,\lambda)$ where $s\in\mathbb{R}^{n}$ is a vector of unit norm, $P$ is a projection of $\mathbb{R}^{n}$, i.e., $P\in\{0,1\}^{n\times n}$ with $P^{2}=P$, and $\varphi$ is a morphism
$\varphi:\Sigma^{*}\longrightarrow O_{n},$ (4)
of the free monoid $\Sigma^{*}$ into the group $O_{n}$ of orthogonal $n\times n$-matrices in $\mathbb{R}^{n\times n}$.
The behaviour of ${\cal Q}$ heavily depends on the topological properties of the semigroup of matrices $\varphi(\Sigma^{*})$. This is why, before returning to quantum automata, we first focus our attention on these matrices for their own sake.
2.1 Topology
The following result is needed in the proof of the main theorem. It will be applied in the particular case of orthogonal matrices. Given a subset $E$ of a finite dimensional vector space, we denote by $\mathbf{C}l(E)$ the topological closure for the topology induced by the Euclidean norm. Given a $k$-tuple of matrices $(M_{1},\ldots,M_{k})$, denote by $f$ the $k$-ary product $f(M_{1},\ldots,M_{k})=M_{1}\cdots M_{k}$ and extend the notation to subsets ${\cal A}$ of $k$-tuples of matrices by posing $f({\cal A})=\{f(M_{1},\ldots,M_{k})\mid(M_{1},\ldots,M_{k})\in{\cal A}\}$. The following result will be applied in several instances of this paper. It says that because we are dealing with compact subsets, the two operators of matrix multiplication and the topological closure commute.
Theorem 1
([7], Theorem 1) Let $\cal C$ be a compact subset of matrices and let ${\cal A}\subseteq{\cal C}^{k}$ be a $k$-ary relation. Then we have ${\mathbf{C}l}(f({\cal A}))=f({\mathbf{C}l}({\cal A}))$.
Consequently, if ${\cal A}$ is a binary relation which is a direct product ${\cal A}_{1}\times{\cal A}_{2}$, we have ${\mathbf{C}l}({\cal A}_{1}{\cal A}_{2})=f({\mathbf{C}l}({\cal A}_{1}\times{\cal A}_{2}))$. It is an elementary result of topology that ${\mathbf{C}l}({\cal A}_{1}\times{\cal A}_{2})={\mathbf{C}l}({\cal A}_{1})\times{\mathbf{C}l}({\cal A}_{2})$ holds. Because of ${\mathbf{C}l}({\cal A}_{1}{\cal A}_{2})=f({\mathbf{C}l}({\cal A}_{1}\times{\cal A}_{2}))=f({\mathbf{C}l}({\cal A}_{1})\times{\mathbf{C}l}({\cal A}_{2}))={\mathbf{C}l}({\cal A}_{1})\ {\mathbf{C}l}({\cal A}_{2})$ we have
Corollary 1
The topological closure of the product of two sets of matrices included in a compact subspace is equal to the product of the topological closures of the two sets.
2.2 Algebraic and semialgebraic sets
Let us give first the definition of algebraic set over the field of real numbers (cf. [3, 13, 26]).
Definition 2
A subset ${\cal A}\subseteq\mathbb{R}^{n}$ is algebraic (over the field of real numbers), if ${\cal A}$ is the zero set of an arbitrary subset $\cal P$ of polynomials of $\mathbb{R}[x_{1},\ldots,x_{n}]$, i.e., for every vector $v\in\mathbb{R}^{n}$,
$v\in{\cal A}\ \Longleftrightarrow\ \forall\ p\in{\cal P}:\ p(v)=0.$ (5)
The set ${\cal A}$ is said to be effective algebraic if the polynomials of $\cal P$ can be computed by an effective procedure (algorithm).
Note that by the Hilbert basis Theorem, one may assume that the set $\cal P$ is finite. Even more, since we are dealing with algebraic sets over $\mathbb{R}$, $\cal P$ can be reduced to a singleton since (5) can be assumed to be to the equation
$\sum_{p\in\cal P}p(x)^{2}=0.$
One can check that the family of algebraic sets is closed under finite unions and intersections. However, it is not closed under complement and projection. The following more general class of subsets enjoys extra closure properties and is therefore more robust.
Definition 3
A subset ${\cal A}\subseteq\mathbb{R}^{n}$ is semialgebraic (over the field of real numbers) if it satisfies one of the two equivalent conditions
(i) ${\cal A}$ is the set of vectors satisfying a finite Boolean combination of predicates of the form $p(x_{1},\ldots,x_{n})>0$ where $p\in\mathbb{R}[x_{1},\ldots,x_{n}]$.
(ii) ${\cal A}$ is first-order definable in the theory of the structure whose domain are the reals and whose predicates are of the form $p(x_{1},\ldots,x_{n})>0$ and $p(x_{1},\ldots,x_{n})=0$ with $p\in\mathbb{R}[x_{1},\ldots,x_{n}]$.
The set ${\cal A}$ is said to be effective semialgebraic if the formula defining ${\cal A}$ can be effectively computed.
We now specify such definitions to square matrices.
Definition 4
A set ${\cal A}\subseteq\mathbb{R}^{n\times n}$ of matrices is algebraic (resp., semialgebraic) if considered as a set of vectors of dimension $n^{2}$, it is algebraic (resp., semialgebraic).
We now combine the notions of zero sets and of topology. In the following we rephrase two results of [8] by emphasizing the main features that serve our purpose. Given a subset $E$ of a group, we denote by $\langle E\rangle$ and by $E^{*}$ the subgroup and the submonoid it generates, respectively.
Theorem 2
Let $S\subseteq\mathbb{R}^{n\times n}$ be a set of orthogonal matrices and let $E$ be any subset of $S$ satisfying $\langle S\rangle=\langle E\rangle$. Then we have ${\mathbf{C}l}(S^{*})={\mathbf{C}l}(\langle E\rangle)$. In particular ${\mathbf{C}l}(S^{*})$ is a group.
Proof.
It is known that every compact subsemigroup of a compact group is a subgroup $G$. Now $S^{*}\subseteq\langle E\rangle$ implies $G={\mathbf{C}l}(S^{*})\subseteq{\mathbf{C}l}(\langle E\rangle)$ and $S\subseteq G$ implies ${\mathbf{C}l}(\langle E\rangle)\subseteq G$ and thus ${\mathbf{C}l}(S^{*})={\mathbf{C}l}(\langle E\rangle)$. ∎∎
The main consequence of the next theorem is that the topological closure of a monoid of orthogonal matrices is algebraic ([26], Ch. 3, Sec. 4, Theorem $5$).
Theorem 3
Let $E$ be a set of orthogonal matrices. Then ${\mathbf{C}l}(\langle E\rangle)$ is a subgroup of orthogonal matrices and it is the zero set of all polynomials $p[x_{1,1},\ldots,x_{n,n}]$ satisfying the conditions
$p(I)=0\ \mbox{ and }p(eX)=p(X)\ \mbox{ for all }e\in E$
Furthermore, if the matrices in $E$ have rational coefficients, the above condition may be restricted to polynomials with coefficients in $\mathbb{Q}$.
2.3 Closure properties of semialgebraic sets
The product of two sets of matrices is defined as
${\cal A}_{1}{\cal A}_{2}=\{M_{1}M_{2}\mid M_{1}\in{\cal A}_{1},M_{2}\in{\cal A}_{2}\}.$ (6)
We define the sandwich operation denoted by $\diamond$ whose first operand is a set of pairs of matrices ${\cal A}\subseteq\mathbb{R}^{n\times n}\times\mathbb{R}^{n\times n}$ and the second operand a set of matrices ${\cal B}\subseteq\mathbb{R}^{n\times n}$ by setting
${\cal A}\diamond{\cal B}=\{XYZ\mid(X,Z)\in{\cal A}\mbox{ and }Y\in{\cal B}\}$
The next operation will be used. Given a bijection
$\pi:\{(i,j)\mid i,j\in\{1,\ldots,n\}\}\rightarrow\{(i,j)\mid i,j\in\{1,\ldots,n\}\}$ (7)
and a matrix $M\in\mathbb{R}^{n\times n}$ denote by $\pi(M)$ the matrix $\pi(M)_{i,j}=M_{\pi(i,j)}$. Extend this operation to subsets of matrices ${\cal A}$. Write $\pi({\cal A})$ to denote the set of matrices $\pi(M)$ for all $M\in{\cal A}$.
The last operation is the sum of square matrices $M_{1},\ldots,M_{k}$ whose result is the square block matrix
$M_{1}\oplus\cdots\oplus M_{k}=\left(\begin{array}[]{cccc}M_{1}&0&\cdots&0\\ 0&M_{2}&\cdots&0\\ \vdots&\vdots&\vdots&\vdots\\ 0&0&0&M_{k}\end{array}\right)$ (8)
These notations extend to subsets of matrices in the natural way. Here we assume that all $k$ matrices have the same size $n\times n$. Observe that if the matrices are orthogonal, so is their sum. Such matrices form a subgroup of orthogonal matrices of size $kn\times kn$.
Now we recall some closure properties of the class of effective semialgebraic matrices investigated in [7]. Let us first consider the operation (6). We adopt the following conventions: we write $\exists^{n}X$ to mean that $X$ is a vector of $n$ bound variables. Furthermore, a vector of $n\times n$ variables can be interpreted as an $n\times n$ matrix of variables. Consider two semialgebraic sets of matrices, say ${\cal A}_{1}$ and ${\cal A}_{2}$, respectively defined by first-order formulas $\Phi_{1}(X_{1})$ and $\Phi_{2}(X_{2})$ where $X_{1}$ and $X_{2}$ are two families of $n^{2}$ free variables viewed as two $n\times n$ matrices of variables. Then ${\cal A}_{1}{\cal A}_{2}$ is defined by the following formula where $X$ is a family of $n^{2}$ free variables viewed as an $n\times n$ matrix
$\exists^{n\times n}X_{1}\;\exists^{n\times n}X_{2}:X=X_{1}X_{2}\wedge\Phi_{1}(X_{1})\wedge\Phi_{2}(X_{2})$
and $X=X_{1}X_{2}$ is an abbreviation for the predicate defining $X$. This shows that the class of effective semialgebraic sets of matrices is closed with respect to (6). The verification of the properties stated below is done similarly.
Proposition 1
Let ${\cal A}_{1},{\cal A}_{2}\subseteq\mathbb{R}^{n\times n}$ be two sets of matrices, and let $\pi$ be a one-to-one mapping as in (7).
1) If ${\cal A}_{1}$ and ${\cal A}_{2}$ are algebraic (resp., effective algebraic) so are ${\cal A}_{1}\cup{\cal A}_{2},{\cal A}_{1}\cap{\cal A}_{2}$, and $\pi({\cal A}_{1})$.
2) If ${\cal A}_{1}$ and ${\cal A}_{2}$ are semialgebraic (resp., effective semialgebraic), so are ${\cal A}_{1}\cup{\cal A}_{2}$, ${\cal A}_{1}{\cal A}_{2}$ and $\pi({\cal A}_{1})$.
Proposition 2
Let ${\cal A}_{1}\subseteq\mathbb{R}^{n\times n}\times\mathbb{R}^{n\times n}$ and ${\cal A}_{2}\subseteq\mathbb{R}^{n\times n}$ be semialgebraic (resp., effective semialgebraic). Then ${\cal A}_{1}\diamond{\cal A}_{2}$ is are semialgebraic (resp., effective semialgebraic).
Proposition 3
Let ${\cal A}$ be a semialgebraic (resp., effective semialgebraic) set of $kn\times kn$ matrices of the form (8). The set $\{X_{1}\cdots X_{k}\mid X_{1}\oplus\cdots\oplus X_{k}\in{\cal A}\}$ is semialgebraic (resp., effective semialgebraic).
Proposition 4
If ${\cal A}_{1},\ldots,{\cal A}_{k}\subseteq\mathbb{R}^{n\times n}$ are algebraic (resp., effective algebraic, semialgebraic, effective semialgebraic), then so is the set ${\cal A}_{1}\oplus\cdots\oplus{\cal A}_{k}$.
2.4 Algebraic Irreducible sets
Let us consider the topological space obtained by equipping $\mathbb{R}^{m}$ with the Zariski topology. Given a non-empty subset $\cal Z$ of $\mathbb{R}^{m}$, $\cal Z$ is said to be irreducible if $\cal Z$ cannot be written as the union of two closed, i.e. algebraic, subsets of $\cal Z$, distinct from $\emptyset$ and $\cal Z$ itself. It can be proven that a non-empty algebraic subset $\cal Z$ of $\mathbb{R}^{m}$ is irreducible if and only if its vanishing ideal is a prime ideal (cf [13], Prop. 1.1.12). Irreducible sets will play a role in this theoretical setting (cf Sec. 6.2). To this purpose, we recall some basic results. The next notation is useful.
Notation. The Zariski closure of a set $\cal A$ will be denoted by $\overline{\cal A}$.
Definition 5
Given non-empty algebraic sets ${\cal V}\subseteq\mathbb{R}^{n}$ and ${\cal W}\subseteq\mathbb{R}^{m}$, a map $f:{\cal V}\rightarrow~{}{\cal W}$ is said to be regular444 By following a well-established vocabulary in Automata Theory, the term regular will be also used in Sec. 4 to name the family of subsets of monoids defined by finite automata (or a morphism of algebraic sets), if there exist $p_{1},\dots,p_{m}\in\mathbb{R}[x_{1},.\ldots,x_{n}]$ (where the $x_{i}$ are indeterminates) such that, for all ${\bf x}=(x_{1},\ldots,n_{n})$ in $\cal V$
$f({\bf x})=(p_{1}({\bf x}),\dots,p_{m}({\bf x})).$ (9)
The following fact is well known (cf [13], Remark 1.3.2).
Lemma 1
If $f:{\cal V}\rightarrow{\cal W}$ is a regular map and $\cal V$ is irreducible, then the Zariski closure $\overline{f({\cal V})}$ of ${f({\cal V})}$ is also irreducible.
[…]
Proposition 5
The family of irreducible algebraic sets is closed with respect to the operations (7), (8) and the (finite) Cartesian product of sets.
Proof.
Let $\cal A$ be an irreducible algebraic set. By Proposition 1, with $\pi$ as in (7), $\pi({\cal A})$ is algebraic so $\overline{(\pi({\cal A}))}=\pi({\cal A})$ and the claim comes from Lemma 1 since $\pi$ is a regular map. For the operation (8) and Cartesian product, the fact that, given irreducible algebraic sets ${\cal A}_{1},\ldots,{\cal A}_{k}$, then so is the set ${\cal A}_{1}\oplus\cdots\oplus{\cal A}_{k}$ (resp., ${\cal A}_{1}\times\cdots\times{\cal A}_{k}$) comes from [13], Prop. 1.3.8-(a).∎∎
The proof of the next lemma is almost immediate.
Lemma 2
Let $H_{1},\ldots,H_{m}$ be subsets of an arbitrary monoid M, every one of which contains the identity $1$ of M. Then $(H_{1}\cup\ldots\cup H_{m})^{*}=(H_{1}\cdots H_{m})^{*}$.
Proof.
[…]
We now introduce an useful notation.
Notation. Given a subset   $\cal A$  of the group $O_{m}$ of orthogonal matrices (over $\mathbb{R}$) of order $m\geq 1$, we write ${\cal A}\in({\cal H})$ if
${\cal A}=\bigcup_{i=1}^{k}{\cal A}_{i},$ (10)
where, for every $i=1,\ldots,k$, ${\cal A}_{i}$ is an irreducible, effective algebraic set that contains the identity matrix $I$.
Proposition 6
Let $\cal A$ be a subset of $O_{nk}$ with ${\cal A}\in({\cal H})$ and let $\Psi$ be a regular map
$\Psi:O_{nk}\;\;\displaystyle\mathop{\longrightarrow}\;\;O_{n}.$
If, for every $i=1,\ldots,k$, $\Psi({\cal A}_{i})$ contains $I$, then $\overline{\Psi({\cal A})^{*}}\in({\cal H})$. In particular $\Psi({\cal A)}$ is effective algebraic.
Proof.
Let ${\cal A}=\bigcup_{i=1}^{k}{\cal A}_{i}$ be a set as in (10). We may suppose that $k=2$, i.e. ${\cal A}={\cal A}_{1}\cup{\cal A}_{2}$, since the general case can be treated similarly. Hence $\Psi({\cal A})=\Psi({\cal A}_{1}\cup{\cal A}_{2})$ so, by Lemma 2,
$\Psi({\cal A})^{*}=(\Psi({\cal A}_{1})\Psi({\cal A}_{2}))^{*}.$
Let $H_{1}=\overline{\Psi({\cal A}_{1})\Psi({\cal A}_{2})}$. By Lemma 1, $\overline{\Psi({\cal A}_{i})}$   $(i=1,2),$ is an algebraic irreducible set. Further, by hypothesis, $\overline{\Psi({\cal A}_{i})}$ contains $I$. Following [12], Sec. 3.1, by using Gröbner bases techniques, since ${\cal A}_{i}$ is effective algebraic and $\Psi$ is a regular map, then one can effectively compute $\overline{\Psi({\cal A}_{i})}$.
Now we observe that
$\overline{\overline{{\Psi(\cal A}_{1})}\cdot\overline{\Psi({\cal A}_{2}})}\ =\ \overline{\Psi({\cal A}_{1})\cdot\Psi({\cal A}_{2})}.$ (11)
Indeed, taking the map m$:O_{n}\times O_{n}\rightarrow O_{n}$ defined, for every $X,Y\in O_{n}$, as m$(X,Y):=X\cdot Y$, one has that m is a regular and thus continuous map (w.r.t. $O_{n}$ endowed with the Zariski topology and $O_{n}\times O_{n}$ with the induced product topology). This implies, for arbitrary sets ${\cal X}_{1},{\cal X}_{2},$ that $\overline{\mbox{m}(\overline{{\cal X}_{1}},\overline{{\cal X}_{2}})}=\overline{\mbox{m}({\cal X}_{1},{\cal X}_{2})},$ and thus (11).
Let $H_{2}=\overline{H_{1}\cdot H_{1}}$. Since $I\in H_{1}$ then one has $I\in H_{2}$. By the very same argument used for $H_{1}$ one has that $H_{2}$ is an irreducible effective algebraic set such that $H_{2}=\overline{\Psi({\cal A})^{2}}$.
Finally, let us define recursively the family of sets of $O_{n}$
$\{H_{i}\}_{i\geq 1},$
where $H_{0}=\{I\}$ and, for every $i\geq 1$,
$H_{i}=\overline{H_{i-1}\cdot H_{1}}.$
By proceeding as in the previous cases, one has that, for every $i\geq 3$, $H_{i}$ is an effective algebraic irreducible set such that $H_{i}=\overline{(\Psi({\cal A}_{1})\Psi({\cal A}_{2}))^{i}}$. and containing $I$. Moreover, since, for each $i\geq 1$, $I\in H_{i}$, then one has the following ascending chain
$H_{1}\;\subseteq\;H_{2}\;\subseteq\;\cdots\;H_{i}\;\subseteq\;\cdots$ (12)
which, by the irreducibility of all of its terms, terminates, i.e., there exists some $m\in\mathbb{N}$ such that (cf Remark 1)
[…]
Finally let ${\cal X}=\bigcup_{i\geq 1}\ H_{i}$. By (13), $\overline{\cal X}=H_{1}\cup\cdots\cup H_{m}$. On the other hand, since, for every $i\geq 1$, $H_{i}=\overline{(\Psi({\cal A}_{1})\Psi({\cal A}_{2}))^{i}}$, then it is readily checked that
$\overline{\cal X}=H_{1}\cup\cdots\cup H_{m}=\overline{\Psi({\cal A})^{*}}.$
This proves the claim. ∎∎
Remark 1
[…]
3 Effectiveness issues
We now return to the $(L,{\cal Q})$ Intersection problem as defined in the Introduction. The main bulk of this decidability result is based upon the following proposition that is obtained by proceeding as in [8]. For the sake of completeness, we recall the proof.
Proposition 7
([7], Proposition 1) Let ${\cal Q}$ be a rational quantum automaton. Let $L\subseteq\Sigma^{*}$ be a formal language such that the set ${\mathbf{C}l}(\varphi(L))$ is effective semialgebraic. It is recursively decidable whether or not $L\cap|{\cal Q}_{>}|=\emptyset$ holds.
Proof.
Equivalently we prove the inclusion $L\subseteq|{\cal Q}_{\leq}|$. We run in parallel two semialgorithms. The first one verifies the noninclusion by enumerating the words $w\in L$ and testing if $||s\varphi(w)P||>\lambda$ holds. The second semialgorithm considers an effectively computable polynomial $\phi(X)$ which defines ${\mathbf{C}l}(\varphi(L))$ and verifies whether the sentence
$\Psi\ \equiv\ \forall X:\phi(X)\Rightarrow sXP\leq\lambda$
holds, which can be achieved by Tarski Seidenberg elimination result. If the inclusion $L\subseteq|{\cal Q}_{\leq}|$ holds then the first semialgorithm cannot answer “yes” and the second semialgorithm will answer “yes”. Now assume that the inclusion does not hold. Then the first semialgorithm will eventually answer “yes”, while the second semialgorithm cannot answer “yes” since this would imply that $\Psi$ holds and thus
$\forall X:X\in{\mathbf{C}l}(\varphi(L))\Rightarrow||sXP||\leq\lambda,$
a contradiction. ∎∎
Proposition 7 shows that the decidability of the problem relies upon the computation of ${\mathbf{C}l}(\varphi(L))$. In the sequel, we will investigate the case of context-free languages. In order to deal with this case, we will use two main tools: first, an algorithm by Derksen et al. [12], already mentioned in the Introduction, that computes the Zariski closure of a finitely generated group of matrices over a field; secondly, a suitable combinatorial structuring of a context-free language based upon some special structures named cycles. The last will be presented in the next section.
4 Preliminaries on context-free languages
For the sake of self-containment and in order to fix notation, we just recall few essential terms of context-free languages and regular sets of monoids, that can be found in all introductory textbooks on theoretical computer science (see, for instance, [4, 15, 16]).
Given a monoid $M,$ a subset of $M$ is said to be regular if it is finite or it is obtained from finite sets by using finitely many times the regular operations, i.e., the union of two sets, the product ()($\cdot$)( ⋅ ) of two sets, and the star $(^{*})$, this last operation associating with a set $\cal X$ the submonoid ${\cal X}^{*}$ generated by $\cal X$. It is worth noticing that the concept of regular set can be equivalently given in terms of $M$-automaton. We will recall this equivalence in the appropriate points (cf Lemma 6, Proposition 10).
A context-free grammar $G$ is a quadruple $\langle V,\Sigma,P,S\rangle$ where $\Sigma$ is the alphabet of terminal symbols, $V$ is the set of variables or nonterminal symbols, $P$ is the set of rules or productions, and $S$ is the axiom of the grammar. A word over the alphabet $\Sigma$ is called terminal. As usual, the nonterminal symbols are denoted by uppercase letters as, for instance, $A$, $B$. An arbitrary rule of the grammar is written as $A\rightarrow\alpha$. The derivation relation of $G$ is denoted by $\mathop{\Rightarrow}^{*}$. The language generated by $G$ is the subset of terminal words $\{w\in\Sigma^{*}:S\displaystyle\mathop{\Rightarrow}^{*}w\}$ denoted by $L(G)$.
The family of context-free languages will be denoted by CFL.
Following [4], Sec. VII.5, we say that a context-free language $L$ is of index $k$, if there exist a context-free grammar $G$ with $L=L(G)$ and an integer $k\in\mathbb{N}$ such that the following property holds: for every word $w\in L(G)$, there exists a derivation $\delta=(S\displaystyle\mathop{\Rightarrow}^{*}w)$ such that the number of occurrences of variables in each sentential form of $\delta$ does not exceed $k$. The family of context-free languages of index $k$ will be denoted by $Ind(k)$ and ${\cal L}(IND_{FIN})$ will denote the family of all the languages of finite index, that is, $\bigcup_{k\in\mathbb{N}}Ind(k)$. One immediately sees that the most simple languages in such a family are those of $LIN$, the linear languages, i.e. those generated by grammars where the right hand side $\alpha$ of every rule $A\rightarrow\alpha$, contains at most one occurrence of nonterminal symbols, that is, $\alpha\in\Sigma^{*}\cup\Sigma^{*}V\Sigma^{*}$. A well-known result proven by Salomaa in [29] shows that Dyck languages cannot be generated by finite index grammars, thus implying ${\cal L}(IND_{FIN})\subset$CFL. The importance of the family ${\cal L}(IND_{FIN})$ is due to the fact that, by a well-known result by Baron and Kuich [2] (up to a technical restriction) the unambiguous context-free languages whose characteristic series in commutative variables are rational, are exactly those in ${\cal L}(IND_{FIN})$. For this reason, such languages are also called “quasi-rational” (see Section 4.2).
4.1 Structuring of CFL
We now find convenient to recall a combinatorial structuring of an arbitrary context-free language (see [7], Sec. 3). Such decomposition follows a recursive scheme proposed by Incitti in [20] based upon the concept defined below.
Definition 6
With each nonterminal symbol $A\in V$ associate the subset of $\Sigma^{*}\times\Sigma^{*}$ defined as
$C_{A}=\{(\alpha,\beta)\in\Sigma^{*}\times\Sigma^{*}:A\displaystyle\mathop{\Rightarrow}^{*}\alpha A\beta\}.$
$C_{A}$ will be called the set of cycles of $A$ [^5] According to the terminology of [15], a cycle is a derivation of the form $A\displaystyle\mathop{\Rightarrow}^{*}\alpha A\beta$. With a minor abuse of language, we will call cycle the pair of contexts $(\alpha,\beta)$ also
It is convenient to define the sandwich operation (cf Sec. 2.3) also for languages in the following way. If $C_{A}$ is the set above and $L^{\prime}$ an arbitrary language, we define the subset of $\Sigma^{*}$
$C_{A}\diamond L^{\prime}=\{uwv\mid(u,v)\in{C_{A}}\mbox{ and }w\in{L^{\prime}}\}$
As the proof of the main theorem proceeds by induction on the number of nonterminal symbols, we need to show how to recombine a grammar from simpler ones obtained by choosing an arbitrary non-axiom symbol as the new axiom and by canceling all the rules involving $S$. This is the reason for introducing the next notation.
Definition 7
Let $G=\langle V,\Sigma,P,S\rangle$ be a context-free grammar. Set $V^{\prime}=V\setminus\{S\}$.
For every $A\in V^{\prime}$, define the context-free grammar $G_{A}=\langle V^{\prime},\Sigma,P_{A},A\rangle$ where the set $P_{A}$ consists of all the rules $B\rightarrow\gamma$ of $G$ of the form
$\quad B\in V^{\prime},\quad\gamma\in(V^{\prime}\cup\Sigma)^{*}$
and denote by $L_{A}$ the language of all terminal words generated by the grammar $G_{A}$.
The next definition introduces the language of terminal words obtained in a derivation where $S$ occurs at the start only.
Definition 8
Let $L^{\prime}(G)$ denote the set of all the words of $\Sigma^{*}$ which admit a derivation
$S\Rightarrow\gamma_{1}\Rightarrow\cdots\Rightarrow\gamma_{\ell}\Rightarrow w$ (14)
where, for every $i=1,\ldots,\ell$, $\gamma_{i}\in(V^{\prime}\cup\Sigma)^{*}$.
The language $L^{\prime}(G)$ can be easily expressed in terms of the languages $L_{A}$ for all $A\in V^{\prime}$. Indeed, consider the set of all rules of the grammar $G$ of the form
$S\rightarrow\beta,\quad\beta\in(V^{\prime}\cup\Sigma)^{*}$ (15)
Factorize every such $\beta$ as
$\beta=w_{1}A_{1}w_{2}A_{2}\cdots w_{\ell}A_{\ell}w_{j_{\ell+1}}$ (16)
where $w_{1},\ldots,w_{\ell+1}\in\Sigma^{*}$ and $A_{1},A_{2},\ldots A_{\ell}\in V^{\prime}$. Next lemma easily follows from the previous definitions.
Lemma 3
With the notation of (16), the language $L^{\prime}(G)$ is the finite union of the languages
$w_{1}L_{A_{1}}w_{2}L_{A_{2}}\cdots w_{\ell}L_{A_{\ell}}w_{j_{\ell+1}}$
when $\beta$ ranges over all rules (15).
Proposition 8
With the previous notation $L$ is a finite union of languages of the form $C_{S}\diamond L^{\prime\prime}$ where
$L^{\prime\prime}=w_{1}L_{A_{1}}w_{2}L_{A_{2}}\cdots w_{\ell}L_{A_{\ell}}w_{\ell+1}$
Proof.
In order to prove the inclusion of the right- into left- hand side, it suffices to consider $w=\alpha u\beta,$ with $u\in L^{\prime}(G)$ and $(\alpha,\beta)\in C_{S}.$ One has $S\displaystyle\mathop{\Rightarrow}^{*}u$ and $S\displaystyle\mathop{\Rightarrow}^{*}\alpha S\beta$ so that $S\displaystyle\mathop{\Rightarrow}^{*}\alpha S\beta\mathop{\Rightarrow}^{*}\alpha u\beta,$ and thus $w\in L$.
Let us prove the opposite inclusion. A word $w\in L$ admits a derivation $S\displaystyle\mathop{\Rightarrow}^{*}w$. If the symbol $S$ does not occur in the derivation except at the start of the derivation, then $w\in L^{\prime}(G)$. Otherwise factor this derivation into $S\displaystyle\mathop{\Rightarrow}^{*}\alpha S\beta\mathop{\Rightarrow}^{*}w$ such that $S$ does not occur in the second part of the derivation except in the sentential form $\alpha S\beta$. Reorder the derivation $\alpha S\beta\mathop{\Rightarrow}^{*}w$ into $\alpha S\beta\mathop{\Rightarrow}^{*}\gamma S\delta\mathop{\Rightarrow}^{*}w$ so that $\gamma,\delta\in\Sigma^{*}$. This implies $w=\gamma u\delta$ for some $u\in L^{\prime}(G)$, which, by Lemma 3, completes the proof. ∎∎
4.2 Structuring of ${\cal L}(IND_{FIN})$
We now recall a characterization, by Ginsburg and Spanier [14] (see also Nivat [24]), of languages of ${\cal L}(IND_{FIN})$, that will be used in Section 6.2. To this purpose, we first recall that, given alphabets $\Sigma_{1}$ and $\Sigma_{2}$, and a family $\cal F$ of languages of $\Sigma_{2}^{*}$, a $\cal F$-substitution is a morphism $\theta:\Sigma_{1}^{*}\rightarrow\Pi(\Sigma_{2}^{*})$ from the free monoid $\Sigma_{1}^{*}$ into the multiplicative monoid $\Pi(\Sigma_{2}^{*})$ of subsets of the free monoid $\Sigma_{2}^{*}$ such that
$\forall\ x\in\Sigma_{1}\quad\theta(x)\in\cal F.$
Let us define, recursively, the family $\{Qrt(k)\}_{k\in\mathbb{N}}$ of quasi-rational languages of rank $k$ (or bounded derivation languages of rank $k$) as:
$Qrt(k)=\left\{\begin{array}[]{ll}LIN&\quad\mbox{if}\ k=1,\\ LIN\circ Qrt(k-1)&\quad\mbox{if}\ k>1,\end{array}\right.$ (17)
where $LIN\circ Qrt(k-1)$ denotes the family of all the languages obtained as the images of linear languages, via $Qrt(k-1)$-substitutions, $\ k>1$, i.e.
$\{\theta(L):L\in LIN,\,\quad\forall\;x\in\Sigma_{1}\;\;\;\theta(x)\in Qrt(k-1)\}.$
The following holds ([14], Theorem 4.2, cf also [4], Sec. VII.5, Theorem 5.2).
Theorem 4
For every $k\geq 1$, $Ind(k)=Qrt(k).$
4.2.1 Composition of grammars
For our next use, we will now describe how quasi-rational languages can be characterized in terms of composition of grammars.
For this purpose, we recall the construction of the grammar that generates the image $\theta(L)$ of a context-free language $L$ over the alphabet $\Sigma_{1}$ under a context-free substitution $\theta:\Sigma_{1}^{*}\rightarrow\Pi(\Sigma_{2}^{*})$, i.e. a $\cal F$-substitution, with $\cal F$=CFL.
We follow verbatim the argument of [4], Sec. II.2. Let ${\cal G}_{1}=\langle V_{1},\Sigma_{1},P_{1},S\rangle$ be the grammar that generates $L$ and let
${\cal G}_{2}=\{G_{x}\ :\ x\in\Sigma_{1}\}$
be the family of grammars where, for every $x\in\Sigma_{1}$,
$G_{x}=\langle V_{x},\Sigma_{2},P_{x},S_{x}\rangle$
is the grammar that generates $\theta(x)$. One may assume that $V_{1}\cap V_{x}=\emptyset$, for all $x\in\Sigma_{1}$. Let be the set
${\cal V}_{2}=\{S_{x}:x\in\Sigma_{1}\}$
and let the copy morphism
$\mbox{c}:(V_{1}\cup\Sigma_{1})^{*}\rightarrow(V_{1}\cup{\cal V}_{2})^{*},$ (18)
defined as: $\mbox{c}(z)=z$, if $z\in V_{1}$, and $\mbox{c}(z)=S_{z}$, if $z\in\Sigma_{1}$. With respect to c, consider the grammar $G_{\mbox{c}}$ that generates $\mbox{c}(L)$. Such a grammar is defined as
$G_{\mbox{c}}=\langle V_{1},\{S_{x}:x\in\Sigma_{1}\},P_{\mbox{c}},S\rangle,$
where $P_{\mbox{c}}=\{A\rightarrow\mbox{c}(\alpha):(A\rightarrow\alpha)\in P\}$. Finally, define
$H=\langle W,\Sigma_{2},Q,S\rangle$ (19)
as the grammar where
$W=V_{1}\ \cup\ \bigcup_{x\in\Sigma_{1}}\ V_{x},\quad\quad Q=P_{\mbox{c}}\ \cup\ \bigcup_{x\in\Sigma_{1}}\ P_{x}.$
The grammar $H$ so obtained will be called the composition of ${\cal G}_{1}$ and ${\cal G}_{2}$ and denoted by
$G={\cal G}_{1}\circ{\cal G}_{2}.$ (20)
Remark 2
By a minor abuse of notation, up to the copy morphism (18), one may identify $\Sigma_{1}$ with ${\cal V}_{2}$. In particular ${V}_{1}\cap{\cal V}_{2}=\emptyset.$
The following lemma holds.
Lemma 4
$\theta(L)=L({\cal G})$.
A composition of $k$ families of grammars $({\cal G}_{1},\ldots,{\cal G}_{k}),$ with $k\geq 1$ will be the grammar
$G={\cal G}_{1}\circ{\cal G}_{2}\circ\cdots\circ{\cal G}_{k},$
obtained by iterating $(k-1)$ times the operation (20) of composition of grammars on the families of context-free grammars ${\cal G}_{1},\ldots,{\cal G}_{k}$ (if the operation is well defined). By (17), Theorem 4 can be formulated as follows.
Theorem 5
Let $L$ be an arbitrary language of $Qrt(k),\ k\geq 1$. Then $L$ is generated by a composition of $k$ families of linear grammars.
Example 1
Let $\cal L$ be the language over $\Sigma_{2}=\{a,b\}$ given by
=${\cal L}=\{a^{n}b^{n}:n\in\mathbb{N}\}^{*}.$
Observe that $\cal L$ is the image of the language $\{x^{n}:n\in\mathbb{N}\}$, $\Sigma_{1}=\{x\}$, under the substitution $\theta:\Sigma_{1}^{*}\rightarrow\Pi(\Sigma_{2}^{*})$, with $\theta(x)=\{a^{n}b^{n}:n\in\mathbb{N}\}$. Hence $\cal L$ is context-free of index $2$. Then $\cal L$ is generated by the composition $G={\cal G}_{1}\circ{\cal G}_{2}$ of the linear grammars ${\cal G}_{1}=\langle V_{1},\Sigma_{1},P_{1},S\rangle$ and ${\cal G}_{2}=\langle V_{2},\Sigma_{2},P_{2},\sigma\rangle$ where
$V_{1}=\{S\},\;\Sigma_{1}=\{\sigma\},\;P_{1}=\{S\rightarrow\varepsilon,\;S\rightarrow\sigma S\},$
and
$V_{2}=\{\sigma\},\;\Sigma_{2}=\{a,b\},\;P_{2}=\{\sigma\rightarrow\varepsilon,\;\sigma\rightarrow a\sigma b\}.$
5 Context-free languages and semialgebraic sets
Now we find convenient to recall (fixing some details of the proof) a property (cf [7], Prop. 7), obtained from the decomposition of Section 4.1. This property is relevant in this study since it states, for an arbitrary context-free language $L$, that the closure ${\mathbf{C}l}(\varphi(L))$ of the image of $L$, under the morphism (4) associated with a quantum automaton, is semialgebraic. To this purpose, the following construction is crucial. With an arbitrary nonterminal $A$ of $G$, associate the set $M_{A}$ defined as
$M_{A}=\{\varphi(u_{1})\oplus\varphi(u_{2})^{\mbox{T}}\mid(u_{1},u_{2})\in C_{A}\},$ (21)
where $\varphi(u_{2})^{\mbox{T}}$ denotes the transpose of $\varphi(u_{2})$. $M_{A}$ will be called the monoid of cycles of $A$ since the following holds.
Lemma 5
Let $A$ be a nonterminal symbol of $G$. $M_{A}$ is a monoid and its closure ${\mathbf{C}l}(M_{A})$ is an algebraic group.
Proof.
Observe first that $M_{A}$ is a subsemigroup of the group $O_{n}\oplus O_{n}.$ Indeed, if $\varphi(u_{1})\oplus\varphi(u_{2})^{\mbox{T}}$ and $\varphi(v_{1})\oplus\varphi(v_{2})^{\mbox{T}}$ are in $M_{A}$ then we have
$(\varphi(u_{1})\oplus\varphi(u_{2})^{\mbox{T}})(\varphi(v_{1})\oplus\varphi(v_{2})^{\mbox{T}})=\varphi(u_{1}v_{1})\oplus\varphi(u_{2})^{\mbox{T}}\varphi(v_{2})^{\mbox{T}},$ (22)
where $(u_{1},u_{2}),(v_{1},v_{2})\in C_{A}$. On the other hand, $(u_{1}v_{1},v_{2}u_{2})\in C_{A}$ and the corresponding matrix of $M_{A}$
$\varphi(u_{1}v_{1})\oplus\varphi(v_{2}u_{2})^{\mbox{T}}$ (23)
equals (22) for $\varphi(v_{2}u_{2})^{\mbox{T}}=\varphi(u_{2})^{\mbox{T}}\varphi(v_{2})^{\mbox{T}}.$ Since $(\varepsilon,\varepsilon)\in M_{A}$, then $M_{A}$ is a monoid. The second statement comes from Theorems 2 and 3. ∎∎
Proposition 9
Let $L$ be a context-free language. The set ${\mathbf{C}l}(\varphi(L))$ is semialgebraic. Moreover, if for every $A\in V$, ${\mathbf{C}l}(M_{A})$ is effective algebraic, then ${\mathbf{C}l}(\varphi(L))$ is effective semialgebraic.
Proof.
By Proposition 8, the language $L$ is a finite union of languages of the form $C_{S}\diamond L^{\prime\prime}$ with
$L^{\prime\prime}=w_{1}L_{A_{1}}w_{2}L_{A_{2}}\cdots w_{\ell}L_{A_{\ell}}w_{\ell+1}$ (24)
where, for every $11\leq i\leq\ell+$ 1, $w_{i}\in\Sigma^{*}$ and $A_{i}\in V^{\prime}$. It suffices to show by induction on the number of nonterminal symbols that the subsets
${\mathbf{C}l}(\varphi(C_{S})\diamond\varphi(L^{\prime\prime}))$ (25)
are algebraic. If $\mathop{\mathrm{Card}}\nolimits(V)=1$, i.e., the set of nonterminal symbols is reduced to $S$ then $L$ is reduced to $C_{S}\diamond L^{\prime}(G)$ and $L^{\prime}(G)$ is finite. We may further assume that there is a sole terminal rule $S\rightarrow w$, so that $L^{\prime}(G)=\{w\}$. By Theorem 1 one has [^6] By Proposition 1, if ${\mathbf{C}l}(M_{S})$ is algebraic, then $\{X\oplus Y^{\mbox{T}}:X\oplus Y\in{\mathbf{C}l}(M_{S})\}$ is so
${\mathbf{C}l}(\varphi(L))=\{X\varphi(w)Y^{\mbox{T}}\mid X\oplus Y\oplus\{\varphi(w)\}\in{\mathbf{C}l}(M_{S}\oplus\varphi(w))\}$ (26)
so that
${\mathbf{C}l}(M_{S}\oplus\varphi(w))={\mathbf{C}l}(M_{S})\oplus{\mathbf{C}l}(\varphi(w))={\mathbf{C}l}(M_{S})\oplus\varphi(w)$
which, by Lemma 5 and Proposition 4, is algebraic. By (26) and Proposition 3, we conclude that ${\mathbf{C}l}(\varphi(L))$ is effective semialgebraic. The basis of the induction is thus proved. Now assume $V$ contains more than one nonterminal symbol. By Theorem 1, Corollary 1 and Eq. (24), we get
${\mathbf{C}l}(\varphi(C_{S})\diamond\varphi(L^{\prime\prime}))=\{XZY^{\mbox{T}}\mid X\oplus Y\oplus Z\in{\mathbf{C}l}(M_{S})\oplus{\mathbf{C}l}(\varphi(L^{\prime\prime}))\}=$
$\{XZY^{\mbox{T}}\mid X\oplus Y\oplus Z\in{\mathbf{C}l}(M_{S})\oplus{\mathbf{C}l}(\varphi(w_{1})\varphi(L_{A_{1}})\cdots\varphi(w_{\ell})\varphi(L_{A_{\ell}})\varphi(w_{\ell+1}))\}$
By Corollary 1 we have
$\begin{array}[]{ll}&{\mathbf{C}l}(\varphi(w_{1})\varphi(L_{A_{1}})\cdots\varphi(w_{\ell})\varphi(L_{A_{\ell}})\varphi(w_{\ell+1}))\\ =&\varphi(w_{1}){\mathbf{C}l}(\varphi(L_{A_{1}}))\cdots\varphi(w_{\ell}){\mathbf{C}l}(\varphi(L_{A_{\ell}}))\varphi(w_{\ell+1}))\end{array}$
which shows, via Proposition 1 and by the induction hypothesis, that this subset is semialgebraic. Then its direct sum with ${\mathbf{C}l}(M_{S})$ is semialgebraic. We conclude by applying Proposition 3. The second part of the statement is obtained from the hypothesis by following the same proof. ∎∎
Lemma 6
Assume that $L$ is a context-free linear language and the quantum automaton is rational. Then, for every variable $A$, ${\mathbf{C}l}(M_{A})$ is effective algebraic.
Proof.
[…]
Proposition 9 together with Lemma 6 finally yield.
Corollary 2
([7], Prop. 22) If $L$ is a context-free linear language and the quantum automaton $\cal Q$ is rational, then its closure ${\mathbf{C}l}(\varphi(L))$ is an effective semialgebraic set. As a consequence, it is recursively decidable whether or not $L\cap|{\cal Q}_{>}|=\emptyset$ holds.
We recall that an intermediate class of languages in between the linear and the finite index ones, is that of metalinear (or ultralinear) languages ${\cal L}(Meta_{Lin})$. By definition, a language $L$ is in ${\cal L}(Meta_{Lin})$ if there exist a context-free grammar $G$ with $L=L(G)$ and an integer $k\in\mathbb{N}$ such that, in every derivation of $G$, the number of occurrences of variables in each sentential form does not exceed $k$. The following corollary holds.
Corollary 3
It is recursively decidable whether or not $L\cap|{\cal Q}_{>}|=\emptyset$ holds, if $L\in{\cal L}(Meta_{Lin})$ and $\cal Q$ is a rational quantum automaton.
Proof.
It is known (see [4, 15]) that $L$ is in ${\cal L}(Meta_{Lin})$ if and only if $L$ is a finite union of finite products of context-free linear languages. Thus the claim follows from Corollary 2, the closure properties of Section 2.3 and Proposition 7. ∎∎
6 Main results
We will prove now that the problem is decidable for two families of languages: the languages generated by the class of restricted matrix context-free grammars and, subsequently, the languages generated by context-free monoidal grammars.
6.1 The case of finite index matrix languages
Now we propose a first non trivial extension of the main results of decidability provided in [7]. Such an extension is based upon the machinery of Subsection 5, and some combinatorial arguments developed by Choffrut, Varricchio and the second author in [10] for the study of rational relations.
To this purpose, following [11], we first recall that a matrix context-free grammar (which we will henceforth simply call matrix grammar) is a tuple $G=\langle V,\Sigma,M,S\rangle$, where $V$, $\Sigma$, and $M$ are respectively the finite sets of nonterminals, terminals, and matrix rules, and $S\in V$ is the start symbol. Each matrix rule $m\in M$ is a finite sequence $m=(p_{1},\ldots,p_{s})$ where each $p_{i}$, $1\leq i\leq s$, is a context-free production from $V$ to $(V\cup\Sigma)^{*}$. For $m\in M$, we define $x\mathrel{\mathop{\Rightarrow}_{m}}y,x,y\in V^{*}$ if $x=x_{0}\mathrel{\mathop{\Rightarrow}_{p_{1}}}x_{1}\mathrel{\mathop{\Rightarrow}_{p_{2}}}\cdots\mathrel{\mathop{\Rightarrow}_{p_{s}}}x_{s}=y$, where $\mathrel{\mathop{\Rightarrow}_{p_{i}}}$ is the standard context-free derivation relation. Let $M^{*}$ be the set of the finite sequences of matrix rules and $\alpha\in M^{*}$. If $\alpha=\varepsilon$, then $x\Rightarrow_{\alpha}x,x\in V^{*}$; if one has
$\alpha=m_{1}m_{2}\cdots m_{s}\mbox{\quad and\quad}x_{0}\mathrel{\mathop{\Rightarrow}_{m_{1}}}x_{1}\mathrel{\mathop{\Rightarrow}_{m_{2}}}\cdots\ x_{s-1}\mathrel{\mathop{\Rightarrow}_{m_{s}}}x_{s},$
with $x_{i}\in V^{*}$, $m_{j}\in M$, $1\leq j\leq s$, $0\leq i\leq s$, then we write $x_{0}\mathrel{\mathop{\Rightarrow}_{\alpha}}x_{s}$. The language generated by $G$, $L(G)=\{w\in\Sigma^{*}\mid S\mathrel{\mathop{\Rightarrow}_{\alpha}}w\}$. The following definition is instrumental.
Definition 9
A matrix grammar $G=(V,\Sigma,M,S)$ is said to be restricted linear of index $k$, or simply restricted, if the set of nonterminals of $G$ has the form $V=\{S\}\cup V^{\prime}$, where $V^{\prime}=V_{1}\cup\dots\cup V_{k},$ and $V_{i}\cap V_{j}=\emptyset,$ for all $i\neq j$, and a matrix rule of $G$ can be only of one of the following forms
  • $(X_{1}\rightarrow u_{1}Y_{1}v_{1},\ldots,X_{k}\rightarrow u_{k}Y_{k}v_{k})$, for some $X_{i},Y_{i}\in V_{i},\ u_{i},\ v_{i}\in\Sigma^{*}$, with $1\leq i\leq k$.
  • For every $X_{i}\in V_{i}$ and $1\leq i\leq k$, $(X_{1}\rightarrow\varepsilon,\ldots,X_{k}\rightarrow\varepsilon)$.
  • $(S\rightarrow X_{1}\cdots X_{k})$, for some $X_{i}\in V_{i}$, with $1\leq i\leq k$.
The main result of this section is the following.
Proposition 10
Let $L$ be a language generated by a restricted grammar. Then one has:
  1. i.
    ${\mathbf{C}l}(\varphi(L))$ is semialgebraic.
  2. ii.
    If the quantum automaton ${\cal Q}$ is rational, then ${\mathbf{C}l}(\varphi(L))$ is effective semialgebraic and the $(L,Q)$ Intersection problem is decidable.
Proof.
(i) Set ${\cal X}=(V^{\prime})^{k}$ and ${\cal Y}=\{X_{1}\cdots X_{k}\in{\cal X}:\exists\ m\in M:m=(S\rightarrow X_{1}\cdots X_{k})\}$. It is easily checked that, given $w\in\Sigma^{*}$, $w\in L$ if and only if there exists some $\omega=(u_{1},v_{1},u_{2},v_{2},\ldots,u_{k},v_{k})\in(\Sigma^{*})^{2k},$ such that
$w=u_{1}v_{1}u_{2}v_{2}\;\cdots\;u_{k}v_{k},$
and $\omega$ belongs to the set $\cal L$ defined as
${\cal L}=\bigcup_{p\in{\cal Y},\;q\in\cal X}\;{\cal L}_{pq},$ (27)
where, for $p=X_{1}\cdots X_{k},\ q=Y_{1}\cdots Y_{k}$, ${\cal L}_{pq}$ is the subset of $(\Sigma^{*})^{2k}$ defined as
${\cal L}_{pq}=\{\;(u_{1},v_{1},\ldots,u_{k},v_{k})\;:\;\exists\ \alpha\in M^{*}:\ X_{1}\cdots X_{k}\Rightarrow_{\alpha}u_{1}Y_{1}v_{1}\cdots u_{k}Y_{k}v_{k}\;\}.$
Let $s=(q_{1},\ldots,q_{n})$ be a repetition-free sequence of elements of $\cal X$, ie, for every positive integers $i,j$, if $i\neq j$, then $q_{i}\neq q_{j}$. Clearly, $n$ is less than or equal to the number of states of $\cal X$ and therefore the set $\cal S$ of all repetition-free sequences is finite. Further, for every $p,q\in\cal X$, let ${\cal S}_{pq}$ be the set
${\cal S}_{pq}=\{s\in{\cal S}\;\mid\;s=(q_{1},\ldots,q_{n}),\;q_{1}=p,\;q_{n}=q\}.$
Now, for every $p,q\in\cal X$, define the subset $E_{pq}$ of $(\Sigma^{*})^{2k}$ as:
$E_{pq}=\{\;(u_{1},v_{1},\ldots,u_{k},v_{k})\;:\;\exists\ m\in M:\ X_{1}\cdots X_{k}\Rightarrow_{m}u_{1}Y_{1}v_{1}\cdots u_{k}Y_{k}v_{k}\;\},$
and the subset $C_{q}$ of $(\Sigma^{*})^{2k}$ as:
$C_{q}=\{\;(u_{1},v_{1},\ldots,u_{k},v_{k})\;:\;\exists\ \alpha\in M^{*}:\ X_{1}\cdots X_{k}\Rightarrow_{\alpha}u_{1}Y_{1}v_{1}\cdots u_{k}Y_{k}v_{k}\;\}.$
One easily checks that $C_{q}$ is a submonoid of the monoid $(\Sigma^{*})^{2k}$. For every $s=(q_{1},\ldots,q_{n})\in\cal S$, define the (possibly empty) subset ${\cal L}_{s}$ of $(\Sigma^{*})^{2k}$ as:
${\cal L}_{s}=C_{q_{1}}E_{q_{1}q_{2}}C_{q_{2}}\cdots C_{q_{n-1}}E_{q_{n-1}q_{n}}C_{q_{n}}$ (28)
By using the very same argument of [10], Theorem 2, one can prove that
${\cal L}_{pq}=\bigcup_{s\in{\cal S}_{pq}}\ {\cal L}_{s}$ (29)
Let now consider the morphism
$\widehat{\varphi}:(\Sigma^{*})^{2k}\rightarrow O_{2nk},$
which maps every element $\omega=(u_{1},v_{1},\ldots,u_{k},v_{k})$ into the orthogonal matrix
$\widehat{\varphi}(\omega)=\varphi(u_{1})\ \oplus\ \varphi(v_{1})\ \oplus\ \cdots\ \oplus\ \varphi(u_{k})\ \oplus\ \varphi(v_{k}),$
of the $2k$-fold product $O_{2nk}$ of $O_{n}.$
Now we prove that ${\mathbf{C}l}({\varphi}(L))$ is semialgebraic. By Proposition 3, it is enough to show that ${\mathbf{C}l}(\widehat{\varphi}(\cal L))$ is so. To this purpose, with an arbitrary nonterminal $q\in{\cal X}$ of $G$, associate the subset $M_{q}$ of $O_{2nk}$ defined as
$\{\varphi(u_{1})\oplus\varphi(v_{1})^{\mbox{T}}\oplus\cdots\oplus\varphi(u_{k})\ \oplus\ \varphi(v_{k})^{\mbox{T}}\mid(u_{1},v_{1},\ldots,u_{k},v_{k})\in C_{q}\}$ (30)
By using the very same argument of Lemma 5, one proves that $M_{q}$ is a monoid and its closure ${\mathbf{C}l}(M_{q})$ is an algebraic group. From this, by Proposition 1, it follows that ${\mathbf{C}l}(\widehat{\varphi}({C}_{q}))$ is algebraic as well. This implies, via Corollary 1 and Proposition 3, that every set ${\mathbf{C}l}(\widehat{\varphi}({\cal L}_{s}))$ of (28) is semialgebraic. Then the semialgebraicity of ${\mathbf{C}l}(\widehat{\varphi}(\cal L))$ follows, by applying Proposition 1, from the last condition and (27), (29).
(ii) Assume that the quantum automaton ${\cal Q}$ is rational. First we prove that ${\mathbf{C}l}(M_{q}))$ is effective algebraic. One proceeds as in the proof of Lemma 6. Indeed, since $L$ is generated by a restricted matrix grammar, then $M_{q}$ is a regular submonoid of $O_{2nk}$. Precisely, it is recognized by the finite $O_{2nk}$–automaton whose states are the elements of $\cal X$, the initial and final states coincide with $q$, and the transitions are of the form
[…]
where, being $s=X_{1}\cdots X_{k},\ t=Y_{1}\cdots Y_{k}$, there exists $m\in M$ such that
$X_{1}\cdots X_{k}\Rightarrow_{m}u_{1}Y_{1}v_{1}\cdots u_{k}Y_{k}v_{k},$
with $\ell(e)\in O_{2nk}$ defined as
$\ell(e)=\varphi(u_{1})\oplus\varphi(v_{1})^{\mbox{T}}\oplus\cdots\oplus\varphi(u_{k})\ \oplus\ \varphi(v_{k})^{\mbox{T}}.$
Hence $M_{q}$ is regular and the group $\langle M_{q}\rangle$ has an effective finite generating set. By applying to $\langle M_{q}\rangle$ the algorithm of Derksen et al. [12], one gets an effective presentation of ${\mathbf{C}l}(M_{q})$. From this, by Proposition 1, it follows that ${\mathbf{C}l}(\widehat{\varphi}({C}_{q}))$ is effective algebraic as well. Finally, since all the steps of the proof (i) are effective, the fact that ${\mathbf{C}l}(\widehat{\varphi}({C}_{q}))$ is effective algebraic implies that ${\mathbf{C}l}({\varphi}(L))$ is so. The claim follows by applying Proposition 7. ∎∎
We now discuss some straightforward corollaries of Proposition 10. First, remark that, by applying the proposition to the matrix grammars of index k=1, we obtains [7], Prop. 22 (cf Corollary 2).
We recall that a bounded semilinear language is a finite union of linear languages which are languages of the form
$L=\{u_{1}^{n_{1}}\cdots u_{k}^{n_{k}}\mid(n_{1},\ldots,n_{k})\in{R}\}$ (31)
for some fixed words $u_{i}\in\Sigma^{*}$ for $i=1,\ldots,k$ and $R\subseteq\mathbb{N}^{k}$ is a linear set, i.e., there exists ${\bf v}_{0},{\bf v}_{1},\ldots,{\bf v}_{\ell}\in\mathbb{N}^{k}$ such that
$R=\{{\bf v}_{0}+\lambda_{1}{\bf v}_{1}+\cdots+\lambda_{\ell}{\bf v}_{\ell}\mid\lambda_{1},\ldots,\lambda_{\ell}\in\mathbb{N}\}$
Lemma 7
Every bounded semi-linear language can be generated by a restricted grammar.
Proof.
It is enough to prove the claim for an arbitrary language $L$ of the form (31). It is checked that $L$ is generated by the restricted grammar $G=\langle V,\Sigma,M,S\rangle$, where $V=\{S,X_{1},\ldots,X_{k}\}$ and $M$ is given by the matrix rules
  • $(S\rightarrow X_{1}\cdots X_{k})$.
  • $(S\rightarrow u_{1}^{{\bf v}_{01}}X_{1}\cdots u_{k}^{{\bf v}_{0k}}X_{k})$, with ${{\bf v}_{0}=({\bf v}_{01}},\ldots,{\bf v}_{0k})$.
  • For every $i=1,\ldots,\ell,$ $(X_{1}\rightarrow u_{1}^{{\bf v}_{i1}}X_{1},\ldots,X_{k}\rightarrow u_{k}^{{\bf v}_{ik}}X_{k})$
[…]
Proposition 10 and Lemma 7 provide a new proof of the following.
Corollary 4
([7], Prop. 21) If $L$ is bounded semilinear then its closure ${\mathbf{C}l}(\varphi(L))$ is semialgebraic. Furthermore, if the quantum automaton ${\cal Q}$ is rational, the $(L,Q)$ Intersection problem is decidable.
We finally observe that Proposition 10 is a non trivial extension of Corollary 2 and Corollary 4. Indeed the following example provides languages, neither bounded nor context-free, defined by restricted grammars.
Example 2
Let $L_{k}$, with $k\in\mathbb{N}$, be the language over $\Sigma=\{a,b\}$ given by $L_{k}=\{u^{k}:u\in\Sigma^{*}\}$. It is easily checked that $L_{k}$ can be generated by the restricted grammar $G=\langle V,\Sigma,M,S\rangle$, where $V=\{S,X_{1},\ldots,X_{k}\}$, and $M$ is given by the matrix rules $(S\rightarrow X_{1}\cdots X_{k})$, $(X_{1}\rightarrow\varepsilon,\ldots,X_{k}\rightarrow\varepsilon)$, and, for every $\sigma\in\Sigma$, $(X_{1}\rightarrow\sigma X_{1},\ldots,X_{k}\rightarrow\sigma X_{k})$.
6.2 The case of monoidal languages
We start with the following definition that is instrumental for this section.
Definition 10
A grammar $G=\langle V,\Sigma,P,S\rangle$ is said to be monoidal of index $k$ if it is the composition
$G={\cal G}_{1}\circ{\cal G}_{2}\circ\cdots\circ{\cal G}_{k},$
of $k$ families of linear grammars where the productions of $G$ satisfy the following conditions:
  • if $\ (X\rightarrow u)\in P,\,\ u\in\Sigma^{*}$ (terminal production)  then $u=\varepsilon$
[…]
A language is said to be monoidal if it is generated by a monoidal grammar of index $k$, for some $k\in\mathbb{N}$. The main result of this section is the following.
Proposition 11
Let $(L,Q)$ be a pair where $Q$ is a rational quantum automaton and $L$ is a language generated by a monoidal grammar $G={\cal G}_{1}\circ{\cal G}_{2}\circ\cdots\circ{\cal G}_{k}$. If the Zariski closures of the monoids of cycles of the grammars of ${\cal G}_{k}$ are irreducible, then ${\mathbf{C}l}(\varphi(L))$ is effective semialgebraic.
We will do the proof of Proposition 11 for $k\leq 2$, i.e., for a monoidal grammar $G={\cal G}_{1}\circ{\cal G}_{2}$ of index $2$, since, by (17), the general case is treated with the same argument. The proof is technically structured in the following lemmata.
Now we find useful to recall a notation introduced in Section 2.4.
Notation. Given a subset $\cal A$ of the group $O_{m}$ of orthogonal matrices (over $\mathbb{R}$), we write
${\cal A}\in({\cal H})$ (32)
if $\cal A$ is a finite union of irreducible, effective algebraic sets, every one of which contains the identity matrix $I$ of $O_{m}$.
Lemma 8
Let $G={\cal G}_{1}$ be a monoidal grammar of index $1$ and let $L=L(G)$. If the Zariski closure of all the monoids of cycles of $G$ are irreducible, effective algebraic sets, then there exist a set $H\in({\cal H})$ and a regular map $\Psi$ such that
${\mathbf{C}l}(\varphi(L))=\Psi(H).$
In particular ${\mathbf{C}l}(\varphi(L))$ is effective semialgebraic and contains $I$.
Proof.
Let $G=\langle V,\Sigma,P,S\rangle$ be a monoidal grammar of index $1$. We rerun the proof of Proposition 9 for $G$. We proceed thus on $\mathop{\mathrm{Card}}\nolimits(V)$. If $\mathop{\mathrm{Card}}\nolimits(V)=1,$ i.e., $V=\{S\}$, then $G$ has a sole monoid of cycles $M_{S}$. Taking into account (26) and the form of the terminal productions of $G$, one has ${\mathbf{C}l}(\varphi(L_{S}))=\Psi({\mathbf{C}l}(M_{S}))$, where $\Psi(X\oplus Y):=XY^{\mbox{T}}.$ From this and the fact that, by hypothesis, ${\mathbf{C}l}(M_{S})$ is irreducible, it follows the claim in this case. The basis of induction is proved.
Let us prove the inductive step. By Proposition 8, and taking into account the form of the productions of $G$, one has
$L_{S}=\bigcup_{i=1}^{m}\,C_{S}\diamond L_{A_{i}},$ (33)
where, for every $i=1,\ldots,m$, $A_{i}\in V^{\prime}=V\setminus\{S\}$. Relation (33) now implies
${\mathbf{C}l}(\varphi(L_{S}))=\bigcup_{i=1}^{m}\,{\mathbf{C}l}(M_{S}\diamond\varphi(L_{A_{i}})).$
By the induction hypothesis, for every $i=1,\ldots,m,$ ${\mathbf{C}l}(\varphi(L_{A_{i}}))=\Psi^{\prime}(H_{i})$, where $H_{i}\in({\cal H})$, $\Psi^{\prime}$ is a regular map (not depending on the variable $A_{i}$) and ${\mathbf{C}l}(\varphi(L_{A_{i}}))$ contains $I$. From this, by a straightforward computation, one gets
${\mathbf{C}l}(\varphi(L_{S}))=\Psi^{\prime\prime}(\,\bigcup_{i=1}^{m}\,{\mathbf{C}l}(M_{S})\oplus H_{i}\,),$ (34)
where
$\Psi^{\prime\prime}(X\oplus Y\oplus Z):=X\Psi^{\prime}(Z)Y^{\mbox{T}}.$
Since ${\mathbf{C}l}(M_{S})$ is effective algebraic and $H_{i}\in({\cal H}),1\leq i\leq m$, then, by Proposition 5, the set $\bigcup_{i=1}^{m}\,{\mathbf{C}l}(M_{S})\oplus H_{i}\in({\cal H})$ as well. Finally the claim comes from (34) taking into account that the map $\Psi^{\prime\prime}$ is regular. ∎∎
Remark 3
Note that the map $\Psi$ of the statement of Lemma 8 does not depend upon the variable used as axiom of $G$.
Let ${\cal G}_{1}=\langle V_{1},\Sigma_{1},P_{1},S\rangle$ be the first grammar of the composition $G={\cal G}_{1}\circ{\cal G}_{2}$. Let $\Sigma_{1}^{*}\times\Sigma_{1}^{*}$ be the multiplicative monoid of binary relations on $\Sigma_{1}^{*}$ and let $\Pi(O_{n}\oplus O_{n})$ be the multiplicative monoid of subsets of $O_{n}\oplus O_{n}$. Define now the morphism between the two monoids
$\zeta:\ \Sigma_{1}^{*}\times\Sigma_{1}^{*}\ \longrightarrow\ \Pi(O_{n}\oplus O_{n})$
as: for every $(u,v)\in\Sigma_{1}^{*}\times\Sigma_{1}^{*}$, with $u=A_{1}\cdots A_{s},\ v=B_{1}\cdots B_{t},\ A_{i},B_{i}\in\Sigma_{1}$
$\zeta(u,v)\ =\ (\varphi(L_{A_{1}})\cdots\varphi(L_{A_{s}})\ \oplus\ \varphi(L_{B_{1}})^{\mbox{T}}\cdots\varphi(L_{B_{t}})^{\mbox{T}})$ (35)
where, for every $i=1,\ldots,t$, $\varphi(L_{B_{i}})^{\mbox{T}}$ denotes the set of matrices
$\{m^{\mbox{T}}\ :\ m\in\varphi(L_{B_{i}})\}.$ (36)
The following lemma holds.
Lemma 9
For every $S\in V_{1}$, there exists an effective computable regular subset $\cal R$ of $\ \Sigma_{1}^{*}\times\Sigma_{1}^{*}$ such that $M_{S}=\zeta({\cal R})$.
Proof.
Let us first define the set
$\widehat{C}_{S}=\{(\alpha,\beta)\in\Sigma_{1}^{*}\times\Sigma_{1}^{*}:(\alpha,\beta^{\sim})\in C_{S}\},$
where $\beta^{\sim}$ denotes the mirror image of $\beta$. Observe that $\widehat{C}_{S}$ is a submonoid of $\Sigma_{1}^{*}\times\Sigma_{1}^{*}$. Moreover one has
$M_{S}=\zeta(\widehat{C}_{S}).$
Indeed, let us check $M_{S}\subseteq\zeta(\widehat{C}_{S}).$ If $\varphi(u)\oplus\varphi(v)^{\mbox{T}}\in M_{S}$, then $S\displaystyle\mathop{\Rightarrow}^{*}uSv$. Since $G={\cal G}_{1}\circ{\cal G}_{2}$, the last derivation can be rewritten as
$S\displaystyle\mathop{\Rightarrow}^{*}\alpha S\beta$ (37)
where (37) is a derivation of ${\cal G}_{1}$ and
$\alpha\displaystyle\mathop{\Rightarrow}^{*}u,\quad\quad\beta\displaystyle\mathop{\Rightarrow}^{*}v,$ (38)
where both the derivations of (38) are in the grammars of ${\cal G}_{2}.$ Now set $\alpha=A_{1}A_{2}\cdots A_{s},$ and $\beta=B_{s}B_{s-1}\cdots B_{1}$, with $A_{i},B_{i}\in\Sigma_{1}\cup\{\varepsilon\},s\geq 0$. From (38), one then has
$u=u_{1}\cdots u_{s},\quad\quad v=v_{s}\cdots v_{1},$
where, for every $i=1,\ldots,s,$ $u_{i}\in L_{A_{i}},$ and $v_{i}\in L_{B_{i}}$. From this, one has
$\varphi(u)\in\varphi(L_{A_{1}})\cdots\varphi(L_{A_{s}}),\quad\quad\varphi(v)^{\mbox{T}}\in\varphi(L_{B_{1}})^{\mbox{T}}\cdots\varphi(L_{B_{t}})^{\mbox{T}}.$ (39)
Finally observe that, by (37), $(\alpha,\beta)\in C_{S}$ so that $(\alpha,\beta^{\sim})\in\widehat{C}_{S}$. By (39) the latter implies $\varphi(u)\oplus\varphi(v)^{\mbox{T}}\in\zeta(\widehat{C}_{S})$. Hence $M_{S}\subseteq\zeta(\widehat{C}_{S})$. The other inclusion is checked similarly.
Since ${\cal G}_{1}$ is a linear grammar, by a well-known construction (cf [4]), one has that $\widehat{C}_{S}$ is an effective computable rational binary relation on $\Sigma_{1}^{*}$. The lemma is proved. ∎∎
Lemma 10
For every $A\in V_{1}$, ${\mathbf{C}l}(M_{A})\in({\cal H}).$ In particular, ${\mathbf{C}l}(M_{A})$ is effective algebraic.
Proof.
By Lemma 9, for every $A\in V_{1}$, there exists an effective computable regular subset $\cal R$ of $\ \Sigma_{1}^{*}\times\Sigma_{1}^{*}$ such that $M_{A}=\zeta({\cal R})$. We will prove the following slightly more general claim:
Claim. For every regular subset $\cal R$ of $\Sigma_{1}^{*}\times\Sigma_{1}^{*}$, ${\mathbf{C}l}(\zeta({\cal R}))$ contains $I$ and ${\mathbf{C}l}(\zeta({\cal R}))=\Psi(H)$ for some $H\in({\cal H})$ (cf (32)) and regular map $\Psi$. Further, if ${\cal R}=R^{*}$, then ${\mathbf{C}l}(\zeta({\cal R}))\in({\cal H})$.
We first show the claim for a finite set $\cal R$. We may suppose $\cal R$ =$=\{r\}$, where $r=(\alpha,\beta)$ with $\alpha=A_{1}A_{2}\cdots A_{s},$ $\beta=B_{1}\cdots B_{t}$, $A_{i},B_{i}\in\Sigma_{1},\ s,t\geq 0$. We assume fist that $\alpha$ and $\beta$ have the same length, i.e., $s=t$. By (35), the fact that $\zeta$ is a monoid morphism and Corollary 1, one gets
${\mathbf{C}l}(\zeta({\cal R}))=\ ({\mathbf{C}l}(\varphi(L_{A_{1}}))\cdots{\mathbf{C}l}(\varphi(L_{A_{s}}))\ \oplus\ {\mathbf{C}l}(\varphi(L_{B_{1}}))^{\mbox{T}}\cdots{\mathbf{C}l}(\varphi(L_{B_{s}}))^{\mbox{T}}),$
taking into account that, for an arbitrary set $\cal A$ of orthogonal matrices, for the set (36), one has ${\mathbf{C}l}({\cal A}^{\mbox{T}})={\mathbf{C}l}({\cal A})^{\mbox{T}}$. By Lemma 8, the previous formula gives
${\mathbf{C}l}(\zeta({\cal R}))=\ (\Psi(H_{1})\cdots\Psi(H_{s})\ \oplus\ \Psi(G_{1})^{\mbox{T}}\cdots\Psi(G_{s})^{\mbox{T}}),$
where $\Psi$ is a regular map and, for every $i=1,\ldots,s$,  $H_{i},G_{i}\in({\cal H})$.
Since each $\Psi(H_{i})$ and $\Psi(G_{i})$ of the last formula, contain $I$ so holds for ${\mathbf{C}l}(\zeta({\cal R}))$. Set $H=H_{1}\oplus\cdots\oplus H_{s}\oplus G_{1}\oplus\cdots\oplus G_{s}$ and observe that, by Proposition 5, $H\in({\cal H})$. Then ${\mathbf{C}l}(\zeta({\cal R}))=\ \Psi^{\prime}(H)$, where $\Psi^{\prime}$ is the regular map defined as
$\Psi^{\prime}(X_{1}\oplus\cdots\oplus X_{s}\oplus Y_{1}\oplus\cdots\oplus Y_{s})=(\Psi(X_{1})\cdots\Psi(X_{s})\oplus\Psi(Y_{1})^{\mbox{T}}\cdots\Psi(Y_{s})^{\mbox{T}}).$
This proves that the claim for $\cal R$ =$=\{r\}$. Finally, if $s>t$, then we can write the word $\beta=B_{1}\cdots B_{t}$ as $\beta=B_{1}\cdots B_{s}$, with $B_{j}=\varepsilon$, for $j=t+1,\ldots,s$. Afterwards, set $L_{B_{j}}=\{\varepsilon\}$ and $\varphi(L_{B_{j}})=\{I\}$, $t+1\leq j\leq s$. Then one proceeds to prove the claim, by repeating the argument above for $s=t$.
We now prove the claim according to the following cases:
[…]
where
${\mathbf{C}l}(\zeta({\cal R}_{1}))=\Psi_{1}(H_{1}),\quad{\mathbf{C}l}(\zeta({\cal R}_{2}))=\Psi_{2}(H_{2}),$ (40)
with $H_{1},H_{2}\in({\cal H})$ and $\Psi_{1},\Psi_{2}$ being regular maps.
(i) By (40) one gets ${\mathbf{C}l}(\zeta({\cal R}))=\Psi(H_{1}\cup H_{2})$, and the claim comes from $H_{1}\cup H_{2}\in({\cal H})$.
(ii) By Corollary 1 and the fact that $\zeta$ is a monoid morphism, (40) yields ${\mathbf{C}l}(\zeta({\cal R}))=\Psi_{1}(H_{1})\Psi_{2}(H_{2}).$ From this, since $I\in\Psi_{1}(H_{1}),\Psi_{2}(H_{2})$ then $I\in{\mathbf{C}l}(\zeta({\cal R}))$. Now, observe that, by Propositions 1, 4, and 5, the condition $H_{1},H_{2}\in({\cal H})$ implies $H_{1}\oplus H_{2}\in({\cal H})$. The claim follows from the fact that
$\Psi_{1}(H_{1})\Psi_{2}(H_{2})=\Psi(H_{1}\oplus H_{2}),$
where $\Psi$ is the regular map defined as $\Psi(X\oplus Y):=\Psi_{1}(X)\Psi_{2}(Y)$.
(iii) Since ${\cal R}=R^{*}$ and the fact that $\zeta$ is a monoid morphism, then one has
${\mathbf{C}l}(\zeta({\cal R}))={\mathbf{C}l}(\zeta({R^{*}}))={\mathbf{C}l}(\zeta({R})^{*}).$
Observe that, from this, it trivially follows that ${\mathbf{C}l}(\zeta({\cal R}))$ contains $I$.
Now taking into account that, by Corollary 1, for an arbitrary set $\cal X$,
${\mathbf{C}l}({\cal X}^{*})={\mathbf{C}l}({\mathbf{C}l}({\cal X})^{*}),$ (41)
one has
${\mathbf{C}l}(\zeta({\cal R}))={\mathbf{C}l}({\mathbf{C}l}(\zeta({R}))^{*}).$
By Theorems 2 and 3, one has
${\mathbf{C}l}({\mathbf{C}l}(\zeta({R}))^{*})=\overline{{\mathbf{C}l}(\zeta({R}))^{*}},$ (42)
where $\overline{{\mathbf{C}l}(\zeta({R}))^{*}}$ is the Zariski closure of ${\mathbf{C}l}(\zeta({R})^{*})$. Recall that, by (40), one has ${\mathbf{C}l}(\zeta({R}))=\Psi(H)$, where $\Psi$ is a regular map and $H\in({\cal H})$. This fact and (42) now implies
${\mathbf{C}l}(\zeta({\cal R}))=\overline{\Psi(H)^{*}}.$
Since $H\in({\cal H})$, we can write $H=\bigcup_{i=1}^{k}\ H_{i}$ as a finite union of irreducible effective algebraic sets $H_{i}$, every one of which containing the identity matrix $I$. This, together with Lemma 2, implies
${\mathbf{C}l}(\zeta({\cal R}))\ =\ \overline{(\Psi(H_{1})\ \cdots\ \Psi(H_{k}))^{*}}.$
Observe that in the formula above, we can write $\Psi(H_{1})\ \cdots\ \Psi(H_{k})$ as
$\widehat{\Psi}(H_{1}\oplus\cdots\oplus H_{k})$, where $\widehat{\Psi}$ is the regular map defined as $\widehat{\Psi}(X_{1}\oplus\cdots\oplus X_{k})=\Psi(X_{1})\cdots\Psi(X_{k})$. Hence we get
${\mathbf{C}l}(\zeta({\cal R}))\ =\overline{(\widehat{\Psi}(H))^{*}},$
where, again, $H=(H_{1}\oplus\cdots\oplus H_{k})\in({\cal H})$ as one can easily checks. The claim follows by applying Proposition 6. This completes the proof. ∎∎
We are now able to prove the main result of this section.
Proof of Proposition 11: As remarked before, it is enough to prove the claim for $k\leq 2$ since, by (17), the general case is treated with the same argument. Let $G=\langle V,\Sigma,P,S\rangle$ and observe that, for every $A\in V$, ${\mathbf{C}l}(M_{A})$ is effective algebraic by hypothesis if $k=1$ and, by Lemma 10, if $k=2$. The claim now follows from Proposition 9. ∎
By Propositions 1 and 7, immediate corollaries of Proposition 11 are the following.
Corollary 5
If the pair $(L,Q)$ satisfies the hypothesis of Proposition 11, then the $(L,Q)$ Intersection problem is decidable.
Corollary 6
Let $(L,Q)$ be a pair where $Q$ is a rational quantum automaton and $L$ is a finite union $L=\bigcup_{i\in{\cal I}}L_{i},$ of monoidal languages, every one of which is such that the corresponding pair $(L_{i},Q)$ satisfies the hypothesis of Proposition 11. Then the $(L,Q)$ Intersection problem is decidable.
An example describing the main steps of Proposition 11 will be given in Section 8.
7 Concluding remarks and open problems
We finally address some questions arising from this study.
[$\alpha$]  One could extend Proposition 11 to a broader class of quantum automata. In this context, the construction of the decidability procedure would require the computation of the Zariski closure of a group $G$ of matrices that is algebraically presented, i.e. $G$ admits a set of generators that is algebraic itself. It may be of interest to investigate conditions assuring the extension of the algorithms of [12, 17, 25] for such groups.
[$\beta$]  It is open whether the $(L,Q)$ Intersection problem is decidable for context-free languages. We conjecture that the answer to this problem is negative. An approach may be to reduce the problem to a so called “birdie problem” [30] (cf also [16], Ch. 8). In such a problem, assuming the feasibility of the computation of the Zariski closure of the groups above, one gets, as a consequence, the decidability of some unsolvable problem.
[$\gamma$]  The $(L,Q)$ Intersection problem is decidable for bounded semi-linear languages. It may be of interest to get extensions of this result for the broader class of languages accepted by the reversal bounded non deterministic counter machines [18]. This extension could exploit a characterization of such languages provided by Ibarra in [19] in terms of some special combinatorial systems called RLGMC-grammars (right-linear grammar with multi-counters). Essentially, a RLGMC-grammar is a regular grammar that associates with each derivation a weight, i.e. a vector in the additive monoid $\mathbb{N}^{d}$. Then a derivation is considered successful if the corresponding weight belongs to the monoid generated by $\bf 1^{d}$ $=(1,1,\ldots,1)$. Here the main bulk is the construction of a formula (see Proposition 7) based on the computation of the Zariski closure of the Intersection of two finitely generated monoids of matrices. The intrinsic difficulty to control the accumulation points of the set – i.e., to exclude those one not coming from the effective functioning of the machines – makes such a computation a non trivial task.
8 Technical details and Examples
We rerun the proofs of Propositions 6 and 11 on the Example 1 of Sec. 4.2.
Example 1 (continued) Let $\cal L$ be the language over the alphabet $\{a,b\}$ given by
=${\cal L}=\{a^{n}b^{n}:n\in\mathbb{N}\}^{*}\ =\ \{a^{n_{1}}b^{n_{1}}\cdots a^{n_{p}}b^{n_{p}}\ :\ n_{1},\ldots,n_{p},\ p\in\mathbb{N}\}.$
Since $\cal L=$ $L^{*},$ with $L=\{a^{n}b^{n}:n\in\mathbb{N}\}$, then
$\mathbf{C}l{(\varphi({\cal L}))}=\mathbf{C}l{(\varphi({L}^{*}))}=\mathbf{C}l{(\varphi({L})^{*})},$
where $\varphi$ is the morphism (4) of the quantum automaton. Recall that $\mathbf{C}l{({\cal A}))}$ and $\overline{{\cal A}}$ denote, respectively, the Euclidean and the Zariski closure of a set $\cal A$. By Theorems 2 and 3, one has $\mathbf{C}l{(\varphi({L})^{*})}=\overline{\varphi({L})^{*}}$. This implies $\mathbf{C}l{(\varphi({\cal L}))}=\overline{\Psi({G})^{*}},$ where
  • $G=M^{*}$ is the monoid generated by the matrix $M\in O_{n}\oplus O_{n}$
    $M=\varphi(a)\oplus\varphi(b)=\left(\begin{array}[]{cc}\varphi(a)&0\\ 0&\varphi(b)\end{array}\right)$
  • $\Psi$ is the regular map defined as $\Psi(A\oplus B):=AB$.
We now prove that, if $\overline{G}$ is irreducible and effective algebraic, then $\overline{\Psi({G})^{*}}$ is in $(\cal H)$ and thus $\mathbf{C}l{(\varphi({\cal L}))}$ is effective algebraic as well.
For this purpose, observe first that $\Psi:O_{n}\oplus O_{n}\rightarrow O_{n}$ is a regular and thus a continuous map. This implies that
$\Psi(\overline{G})\subseteq\overline{\Psi(G)},$
which, in turn, implies
$\overline{\Psi(\overline{G})}\ =\ \overline{\Psi(G)}.$ (43)
Let $H_{1}=\overline{\Psi(\overline{G})}$. By Lemma 1, $H_{1}$ is an algebraic irreducible set. Moreover, being $G$ a group, one has $I\in\Psi(G)$ so $I\in H_{1}$. Since $G$ is finitely generated as a monoid, then, by using Derksen’s algorithm, one can effectively calculate $\overline{G}$. Afterwards, following [12], Sec. 3.1, by using Gröbner bases techniques, since $\overline{G}$ is effective algebraic, then one can effectively compute $H_{1}$.
Let $H_{2}=\overline{H_{1}\cdot H_{1}}$. Observe first that $I\in H_{1}$ implies $I\in H_{2}$. Similarly to the previous case, one observes that, for arbitrary subsets ${\cal A,B}$ of $O_{n}$
[…]
This implies that $\overline{\mbox{m}(\overline{{\cal A}},\overline{{\cal B}})}=\overline{\mbox{m}({\cal A},{\cal B})}.$ By using the very same argument for $H_{1}$ one verifies that $H_{2}$ is an irreducible effective algebraic set. Moreover, by (43) and (44), one has $H_{2}=\overline{\Psi(G)^{2}}$.
Finally, let us define recursively the family $\{H_{i}\}_{i\geq 1}$ of sets of $O_{n}$, where $H_{0}=\{I\}$ and, for every $i\geq 1$, $H_{i}=\overline{H_{i-1}\cdot H_{1}}.$ Getting an ascending chain
$H_{1}\;\subseteq\;H_{2}\;\subseteq\;\cdots\;H_{i}\;\subseteq\;\cdots$
of irreducible, effective algebraic sets, then there exists some $m\in\mathbb{N}$ such that
[…]
Finally let ${\cal X}=\bigcup_{i\geq 1}\ H_{i}$. By (45), $\overline{\cal X}=H_{1}\cup\cdots\cup H_{m}$. On the other hand, since, for every $i\geq 1$, $H_{i}=\overline{\Psi(G)^{i}}$, then it is checked that
$\overline{\cal X}=H_{1}\cup\cdots\cup H_{m}=\overline{\Psi(G)^{*}}.$
Hence $\overline{\Psi(G)^{*}}$ is effective algebraic.