### v3

Abstract
In this paper, we provide some conditions of algebraic and combinatorial flavour that make decidable the following problem: given a measure-once finite quantum automaton and a context-free grammar of finite index or a simple matrix context-free grammar, the set intersection of the language accepted by the automaton and the language generated by the grammar, is empty or not.
Keywords: Quantum automata, Context-free languages, Algebraic groups, Matrix semigroups, Decidability
AMS subject classifications: 81P68, 68Q45, 03D05
1 Introduction
A fundamental issue in the mathematical theory of formal languages concerns the decidability of the Emptiness Problem: given a model of computation $\cal Q$, is the language accepted by $\cal Q$, $L({\cal Q})$, equal to the empty set? A natural generalization of such a problem, called Intersection Problem, arises by relativizing it in the following way: one considers a predefined family $\cal L$ of (effectively defined) languages and asks whether the intersection of $L({\cal Q})$ with a language taken arbitrarily from $\cal L$ is empty. Notice that, if $\cal L$ is the singleton set containing the free monoid of words over the input alphabet of $\cal Q$, one immediately obtains the Emptiness Problem.
In this paper, by continuing an investigation of Bertoni, Choffrut et al. [9, 10], we will study the decidability of the Intersection Problem w.r.t. quantum finite automata as the family $\cal Q$ and context-free languages, together with some other strictly related languages, as the family $\cal L$. Quantum automata were introduced at the beginning of 2000’s as a new model of language recognizer, that can be obtained by imposing the quantum paradigm on classical finite automata. Several models of quantum automata have been proposed in the literature, which essentially differ in what quantum measurements are allowed (see [1, 8, 21] for reference and updated texts). In this paper, we consider the model of a quantum automaton, called “measure-once”, introduced by Moore and Crutchfield in [33]. Such a model is the simplest and one of the most studied in the literature since it provides the description of good-featured quantum devices of “small size”. In this regard, it is worth recalling that a recent contribution of Mereghetti and Palano describes a method for the physical implementation of measure-once quantum automata for the recognition of a well-known family of periodic languages [31, 32].
Numerous publications have investigated the decision properties of quantum automata. The starting point of that investigation is a result of Blondel, Jeandel, Koiran, and Portier who showed in [11] that, given a quantum automaton $\cal Q$ (in the sense of [33]), it is decidable whether the language recognized by $\cal Q$ has a non trivial intersection with the free monoid generated by the input alphabet of $\cal Q$ (the Emptiness Problem, aforementioned). It is worth remarking that these problems are related to some computational issues concerning matrices. Precisely, a crucial step in the construction of the decision procedure for both the Emptiness and Intersection Problem amounts to effectively compute the Zariski closure of a finitely generated group of matrices over $\mathbb{Q}$. A striking result proved by Derksen, Jeandel and Koiran in [15] shows the existence of an appropriate algorithm for that task.
Before coming back to the Intersection Problem, we find then convenient to recall some developments of the research on that aspect. Two results have been obtained recently in connection with the algorithm of [15]. In 2022 Nosan, Pouly, Schmitz, Shirmohammadi, and Worrell, following an alternative approach for such computation, have obtained in [35] a bound on the degree of the polynomials that define the Zariski closure of the set and showed that the latter can be computed in elementary time.
The second was obtained in 2023 by Hrushovski, Ouaknine, Pouly, and Worrell for the study of the effective computation of polynomial invariants for affine programs [23, 22]. Precisely, it is exhibited an algorithm that, given a finite set of square matrices of the same size with coefficients in $\mathbb{Q}$, computes the Zariski closure of the semigroup generated by the set.
It should be pointed out that the decidability issues for matrix semigroups is a challenging problem. Even in small size, a well-known result by Paterson shows that the problem of freeness and that of membership are both undecidable for matrix semigroups over $\mathbb{Z}$ of size $3$ (see [12]). In contrast, in 2017, a surprising result of Potapov and Semukhin [38] shows that the membership is decidable for non-singular matrix integer semigroups of size $2$.
We now come back to the Intersection Problem and give a formal presentation of our work. A finite quantum automaton (in the sense of [33]) is a quadruple
${\cal Q}=(s,\varphi,P,\lambda),$ (1)
where $s\in\mathbb{R}^{n}$ is a row-vector of unit Euclidean norm, $P$ is an orthogonal projection matrix of size ${n}$, $\varphi:\Sigma^{*}\rightarrow O_{n}$ is a morphism from the free monoid $\Sigma^{*}$, generated by the input alphabet $\Sigma$ of $\cal Q$, into the group $O_{n}$ of orthogonal $n\times n$-matrices over $\mathbb{R}$ and the threshold $\lambda$ has value in $\mathbb{R}$.
We are mainly interested in effective properties which require the quantum automaton to be effectively given. We thus consider the model of rational quantum automaton, i.e., where all the coefficients of the components of (1) and $\lambda$ are rational numbers. For a real threshold $\lambda$, the languages recognized by $\cal Q$ with strict and nonstrict threshold $\lambda$ are respectively
$|{\cal Q}_{>}|=\{w\in\Sigma^{*}\mid||s\varphi(w)P||^{2}>\lambda\},\quad|{\cal Q}_{\geq}|=\{w\in\Sigma^{*}\mid||s\varphi(w)P||^{2}\geq\lambda\},$
where |||||$\cdot$||| | ⋅ | | denotes the Euclidean norm of vectors. Let ${\cal L}$ be a family of effectively defined languages. The problem we tackle is the following:
$(L,{\cal Q})$ Intersection
Input: a language $L$ in a family ${\cal L}$ of languages and a rational quantum automaton ${\cal Q}$.
Question: does $L\cap|{\cal Q}_{>}|=\emptyset$ hold?
As already pointed out, the decidability of this problem has been proven in [11] in the specific case $L$ $=\Sigma^{*}$. This result contrasts with the corresponding problems (strict and non strict, respectively) for probabilistic automata. Indeed, for these automata, the Emptiness Problems w.r.t. $|{\cal Q}_{>}|$ and $|{\cal Q}_{\geq}|$ are both undecidable [37], Thm 6.17 (see also [7] for the undecidability of the property to be isolated for a cut point). Also it contrasts with the undecidability of the Emptiness Problem of $|{\cal Q}_{\geq}|$ for quantum automata [11]. It is also proven in [29] that both the problems (strict and non strict, respectively) remain undecidable for the “measure many” model of quantum automata, a model not computationally equivalent to the “measure once”, introduced by Kondacs and Watrous in [30]. In order to describe the main contributions of the paper, it may be useful to recall some lines of the technique used in the proof of the Intersection Problem in [10]. For this purpose, following the same strategy of [11], we first observe that the Intersection Problem is equivalent to the inclusion
$\varphi(L)\subseteq|{\cal Q}_{\leq}|.$ (2)
Since the function
$M\rightarrow||sMP||^{2},$
– where $M$ is an arbitrary matrix in $\mathbb{R}^{n\times n}$, and $s,P$ are components of (1) – is continuous, it is sufficient to prove that, for all matrices $M$ in the Euclidean closure ${\mathbf{C}l}(\varphi(L))$ of $\varphi(L)$, the condition $||sMP||^{2}\leq\lambda$ holds (see Section 3). Such condition is then effectively tested by expressing the property (2) in first-order logic of the field of reals, which, in turn, consists to effectively compute a representation of ${\mathbf{C}l}(\varphi(L))$ in term of a semialgebraic set (a more general property than algebraic, which allows one to cope with the fact that algebraic sets are not closed w.r.t. the operation of product of sets). Afterwards, one applies on the constructed formula the Tarski-Seidenberg quantifier elimination to verify whether (2) holds true or not.
It turns out that, in the case the language $L$ is context-free, the semialgebraicity of ${\mathbf{C}l}(\varphi(L))$, and its effective computation (as a semialgebraic set), are reduced to those of an algebro-combinatorial object that is associated with the grammar $G$ generating $L$, and with every variable $A$ of $G$: the monoid of cycles of $A$. Such monoid, denoted $M_{A}$, corresponds (up to a technical detail) to the matrix image, under taking the morphism $\varphi$ of (1), of the language of cycles in $G$ associated with $A$ (see Eq. (20)).
The reduction process mentioned above is made possible by resorting to two ingredients: the fact that the Euclidean closure ${\mathbf{C}l}(M_{A})$ is an algebraic set over the field $\mathbb{R}$ of real numbers, because of a deep result of Algebraic Geometry (see Theorem 3) and because of a suitably defined effective structuring of the computations, i.e., the derivations, of $G$ (Sections 4.1, 4.2 and 6.1). In this theoretical setting, the work in [10] and this work provide the following picture of the Intersection Problem:
  • If $L$ is a context-free language, then ${\mathbf{C}l}(\varphi(L))$ is semialgebraic. Moreover, if, for every variable $A$ of $G$, ${\mathbf{C}l}(M_{A})$ is effectively computable, then ${\mathbf{C}l}(\varphi(L))$ is so as well (Proposition 9).
  • If, for every variable $A$ of $G$, $M_{A}$ is a regular submonoid, then ${\mathbf{C}l}(\varphi(L))$ is effectively computable. This is guaranteed by applying Proposition 9 and the results [15, 23, 35] mentioned above. This, in particular, covers the cases of context-free linear languages, bounded semi-linear languages and a new class of languages, proposed in this paper, generated by restricted matrix context-free grammars which non trivially extends the previous ones (Proposition 10).
  • The computation of ${\mathbf{C}l}(\varphi(L))$ is achieved for a subfamily of finite index context-free languages, called monoidal, by considering the property of (algebraic) irreducibility on the Zariski closure over $\mathbb{R}$ of a family of monoids of cycles canonically associated with $G$ (Proposition 11).
The new results of the paper are the following. In Section 6.1, we extend the main contribution of [10] to a subfamily of simple matrix languages. Simple matrix languages have been introduced by Ibarra in [24] as a proper class of the deterministic linear-bounded acceptor languages, defined by a restricted form of matrix context-free grammars (see [27] for recent developments). They contain, in particular, context-free languages and inherit several structural and decidable properties. Precisely, we prove that, whenever the quantum automaton is rational, such extension (Proposition 10) allows one, on one hand, to recover both the previous two cases – i.e., the case of linear context-free languages and that of bounded semi-linear languages – as immediate corollaries. On the other hand, it shows the decidability of the problem for languages of exponential growth that are not context-free. In the proof, a suitable decomposition of the derivations in these grammars is used. Such a decomposition is based upon some combinatorial tools proposed by Choffrut, Varricchio and the second author in [13] for the study of the counting functions of rational relations. In particular, this decomposition permits a refinement of the machinery developed in [10] for the manipulation of the monoids of cycles of these grammars.
In Section 6.2, we investigate the Intersection Problem for languages defined by finite index context-free grammars. These are grammars $G$ where each word $w$ in the language generated by $G$ is obtained by some derivation $\delta$, whose index is uniformly bounded, i.e., the number of variables in each sentential form of $\delta$ is bounded by an integer not depending on $w$. A result by Ginsburg and Spanier [18] (see also Nivat [34]) provides a characterization of such languages in terms of composition of grammars: precisely, each language of this type is generated by a grammar $G$ given by the composition
$G={\cal G}_{1}\circ{\cal G}_{2}\circ\cdots\circ{\cal G}_{k},$ (3)
[…]
We prove that, regardless of the length $k$ of the composition (3) for $G$, if the Zariski closure over $\mathbb{R}$ of the monoids of cycles of the grammars of the lowest level ${\cal G}_{k}$ of (3), are (algebraic) irreducible, then one can compute the Euclidean closure ${\mathbf{C}l}(\varphi(L))$ of $\varphi(L)$ (Proposition 11), thus making the Intersection Problem decidable (Corollaries 5 and 6).
Extensions of such result to broader families of finite index context-free languages seem related to possible extensions of the algorithms [15, 23, 35] (cf Section 7, [$\alpha$]). The general case of context-free languages is open and it could be even possible that the problem is undecidable (cf Section 7, [$\beta$]).
The paper is structured in the following way. In Section 2 preliminaries on semialgebraic and irreducible algebraic sets are presented. In Section 3 we recall an effective condition to which the decidability of the problem under discussion can be reduced. Section 4 provides combinatorial structurings of context-free languages in the general case and in the finite index case. Section 5 recalls some results relating context-free grammars and semialgebraic sets. In Section 6 the main results of the paper are presented. Section 7 contains some concluding remarks and open problems. In the Appendix, the construct of composition of grammars is given in detail.
The results of this paper have been announced at RP 2024 [5]. The present paper contains fully detailed proofs and examples of applications and construction.
2 Preliminaries on semialgebraic sets
Let us give first the definition of quantum automaton (cf. [9, 11, 33]).
Definition 1
A quantum automaton ${\cal Q}$ is a quadruple $(s,\varphi,P,\lambda)$ where $s\in\mathbb{R}^{n}$ is a vector of unit norm, $P$ is an orthogonal matrix of size $n$ with $P^{2}=P$, and $\varphi$ is a morphism
$\varphi:\Sigma^{*}\longrightarrow O_{n},$ (4)
of the free monoid $\Sigma^{*}$ into the group $O_{n}$ of orthogonal $n\times n$-matrices in $\mathbb{R}^{n\times n}$.
Remark 1
Throughout the paper we will assume that the linear representation of every quantum automaton (with possibly complex entries) is given, as in Definition 1, by matrices over the field of real numbers. This is a merely technical restriction since any quantum automaton can be simulated by another one of the latter type. We refer the reader to the Section “Complex versus real entries” of the paper [11], where this issue is very clearly and thoroughly explained.
The behaviour of ${\cal Q}$ heavily depends on the topological properties of the semigroup of matrices $\varphi(\Sigma^{*})$. This is why, before returning to quantum automata, we first focus our attention on these matrices for their own sake.
2.1 Topology
The following result is needed in the proof of the main theorem. It will be applied in the particular case of orthogonal matrices. Given a subset $E$ of a finite dimensional vector space, we denote by $\mathbf{C}l(E)$ the topological closure for the topology induced by the Euclidean norm. Given a $k$-tuple of matrices $(M_{1},\ldots,M_{k})$, denote by $f$ the $k$-ary product $f(M_{1},\ldots,M_{k})=M_{1}\cdots M_{k}$ and extend the notation to subsets ${\cal A}$ of $k$-tuples of matrices by posing $f({\cal A})=\{f(M_{1},\ldots,M_{k})\mid(M_{1},\ldots,M_{k})\in{\cal A}\}$. The following result will be applied in several instances of this paper. It says that because we are dealing with compact subsets, the two operators of matrix multiplication and the topological closure commute.
Theorem 1
([10], Theorem 1) Let $\cal C$ be a compact subset of matrices and let ${\cal A}\subseteq{\cal C}^{k}$ be a $k$-ary relation. Then we have ${\mathbf{C}l}(f({\cal A}))=f({\mathbf{C}l}({\cal A}))$.
Thus, if ${\cal A}$ is a binary relation which is a direct product ${\cal A}_{1}\times{\cal A}_{2}$, we have ${\mathbf{C}l}({\cal A}_{1}{\cal A}_{2})=f({\mathbf{C}l}({\cal A}_{1}\times{\cal A}_{2}))$. It is an elementary result of topology that ${\mathbf{C}l}({\cal A}_{1}\times{\cal A}_{2})={\mathbf{C}l}({\cal A}_{1})\times{\mathbf{C}l}({\cal A}_{2})$ holds. Because of ${\mathbf{C}l}({\cal A}_{1}{\cal A}_{2})=f({\mathbf{C}l}({\cal A}_{1}\times{\cal A}_{2}))=f({\mathbf{C}l}({\cal A}_{1})\times{\mathbf{C}l}({\cal A}_{2}))={\mathbf{C}l}({\cal A}_{1})\ {\mathbf{C}l}({\cal A}_{2})$, we have the following
Corollary 1
The topological closure of the product of two sets of matrices included in a compact subspace is equal to the product of the topological closures of the two sets.
2.2 Algebraic and semialgebraic sets
Let us give first the definition of algebraic set over the field of real numbers (cf. [4, 17, 36]).
Definition 2
A subset ${\cal A}\subseteq\mathbb{R}^{n}$ is algebraic (over the field of real numbers), if ${\cal A}$ is the zero set of an arbitrary subset $\cal P$ of polynomials of $\mathbb{R}[x_{1},\ldots,x_{n}]$, i.e., for every vector ${\bf v}\in\mathbb{R}^{n}$,
${\bf v}\in{\cal A}\ \Longleftrightarrow\ \forall\ p\in{\cal P}:\ p({\bf v})=0.$ (5)
Note that by the Hilbert’s basis theorem (see [17], Ch. 1, Sec. 1, Theorem $1$), one may assume that the set $\cal P$ is finite. Even more, since we are dealing with algebraic sets over $\mathbb{R}$, $\cal P$ can be reduced to a singleton since (5) can be assumed to be to the equation
$\sum_{p\in\cal P}p({\bf v})^{2}=0.$
One can check that the family of algebraic sets is closed under finite unions and intersections. However, it is not closed under complement and projection. The following more general class of subsets enjoys extra closure properties and is thus more robust. The equivalence of the two definitions below is guaranteed by Tarski-Seidenberg quantifier elimination result (cf [4], Ch. 2, Sec. 2.3 and Sec. 2.5, Cor. 2.95).
Definition 3
A subset ${\cal A}\subseteq\mathbb{R}^{n}$ is semialgebraic (over the field of real numbers) if it satisfies one of the two equivalent conditions
(i) ${\cal A}$ is the set of vectors satisfying a finite Boolean combination of predicates of the form $p(x_{1},\ldots,x_{n})>0$ where $p\in\mathbb{R}[x_{1},\ldots,x_{n}]$.
(ii) ${\cal A}$ is first-order definable in the theory of the structure whose domain are the reals and whose predicates are of the form $p(x_{1},\ldots,x_{n})>0$ or $p(x_{1},\ldots,x_{n})=0$ with $p\in\mathbb{R}[x_{1},\ldots,x_{n}]$.
We now instantiate the definitions above to the setting of square matrices.
Definition 4
A set ${\cal A}\subseteq\mathbb{R}^{n\times n}$ of matrices is algebraic (resp., semialgebraic) if considered as a set of vectors of dimension $n^{2}$, it is algebraic (resp., semialgebraic).
Notational convention. In connection with the $(L,{\cal Q})$ Intersection Problem, we will adopt the following terminology. A set ${\cal A}$ of matrices will be called effective algebraic (resp., effective semialgebraic) if the polynomials (resp., the formula) defining ${\cal A}$ can be algorithmically computed from the input of the problem, i.e., an effectively defined language $L$ and a finite quantum automaton $\cal Q$.
We now combine the notions of zero sets and of topology. In the following we rephrase two results of [11] by emphasizing the main features that serve our purpose. Given a subset $E$ of a group, we denote by $\langle E\rangle$ and by $E^{*}$ the subgroup and the submonoid generated by $E$, respectively.
Theorem 2
Let $S\subseteq\mathbb{R}^{n\times n}$ be a set of orthogonal matrices and let $E$ be any subset of $S$ satisfying $\langle S\rangle=\langle E\rangle$. Then we have ${\mathbf{C}l}(S^{*})={\mathbf{C}l}(\langle E\rangle)$. In particular ${\mathbf{C}l}(S^{*})$ is a group.
Proof.
It is known that every compact subsemigroup of a compact group is a subgroup $G$. Now $S^{*}\subseteq\langle E\rangle$ implies $G={\mathbf{C}l}(S^{*})\subseteq{\mathbf{C}l}(\langle E\rangle)$ and $S\subseteq G$ implies ${\mathbf{C}l}(\langle E\rangle)\subseteq G$ and thus ${\mathbf{C}l}(S^{*})={\mathbf{C}l}(\langle E\rangle)$. ∎∎
The main consequence of the next theorem is that the topological closure of a monoid of orthogonal matrices is algebraic ([36], Ch. 3, Sec. 4, Theorem $5$, [10], Theorem $7$).
Theorem 3
Let $E$ be a set of orthogonal matrices. Then ${\mathbf{C}l}(\langle E\rangle)$ is a subgroup of orthogonal matrices and it is the zero set of all polynomials $p[x_{1,1},\ldots,x_{n,n}]$ satisfying the conditions
$p(I)=0\ \mbox{ and }p(eX)=p(X)\ \mbox{ for all }e\in E$
Furthermore, if the matrices in $E$ have rational coefficients, the above condition may be restricted to polynomials with coefficients in $\mathbb{Q}$.
2.3 Closure properties of semialgebraic sets
The product of two sets of matrices is defined as
${\cal A}_{1}{\cal A}_{2}=\{M_{1}M_{2}\mid M_{1}\in{\cal A}_{1},M_{2}\in{\cal A}_{2}\}.$ (6)
We define the sandwich operation denoted by $\diamond$ whose first operand is a set of pairs of matrices ${\cal A}\subseteq\mathbb{R}^{n\times n}\times\mathbb{R}^{n\times n}$ and the second operand a set of matrices ${\cal B}\subseteq\mathbb{R}^{n\times n}$ by setting
${\cal A}\diamond{\cal B}=\{XYZ\mid(X,Z)\in{\cal A}\mbox{ and }Y\in{\cal B}\}$
Given a bijection
$\pi:\{(i,j)\mid i,j\in\{1,\ldots,n\}\}\rightarrow\{(i,j)\mid i,j\in\{1,\ldots,n\}\}$ (7)
and a matrix $M\in\mathbb{R}^{n\times n}$ we define a new operation by denoting $\pi(M)$ as the matrix $\pi(M)_{i,j}=M_{\pi(i,j)}$. Extend this operation to subsets of matrices ${\cal A}$. Write $\pi({\cal A})$ to denote the set of matrices $\pi(M)$ for all $M\in{\cal A}$.
The last operation is the sum of square matrices $M_{1},\ldots,M_{k}$ whose result is the square block matrix
$M_{1}\oplus\cdots\oplus M_{k}=\left(\begin{array}[]{cccc}M_{1}&0&\cdots&0\\ 0&M_{2}&\cdots&0\\ \vdots&\vdots&\vdots&\vdots\\ 0&0&0&M_{k}\end{array}\right)$ (8)
These notations extend to subsets of matrices in the natural way. Here we assume that all $k$ matrices have the same size $n\times n$. Observe that if the matrices are orthogonal, so is their sum. Such matrices form a subgroup of orthogonal matrices of size $kn\times kn$.
Now we recall some closure properties of the class of semialgebraic sets of matrices (see [10], Sec. 2.4).
Proposition 1
Let ${\cal A}_{1},{\cal A}_{2}\subseteq\mathbb{R}^{n\times n}$ be two sets of matrices, and let $\pi$ be a one-to-one mapping as in (7).
1) If ${\cal A}_{1}$ and ${\cal A}_{2}$ are algebraic (resp., effective algebraic) so are ${\cal A}_{1}\cup{\cal A}_{2},{\cal A}_{1}\cap{\cal A}_{2}$, and $\pi({\cal A}_{1})$.
2) If ${\cal A}_{1}$ and ${\cal A}_{2}$ are semialgebraic (resp., effective semialgebraic), so are ${\cal A}_{1}\cup{\cal A}_{2}$, ${\cal A}_{1}{\cal A}_{2}$ and $\pi({\cal A}_{1})$.
Proposition 2
Let ${\cal A}_{1}\subseteq\mathbb{R}^{n\times n}\times\mathbb{R}^{n\times n}$ and ${\cal A}_{2}\subseteq\mathbb{R}^{n\times n}$ be semialgebraic (resp., effective semialgebraic). Then ${\cal A}_{1}\diamond{\cal A}_{2}$ is semialgebraic (resp., effective semialgebraic).
Proposition 3
If ${\cal A}_{1},\ldots,{\cal A}_{k}\subseteq\mathbb{R}^{n\times n}$ are algebraic (resp., effective algebraic, semialgebraic, effective semialgebraic), then so is the set ${\cal A}_{1}\oplus\cdots\oplus{\cal A}_{k}$.
Proposition 4
Let ${\cal A}$ be a semialgebraic (resp., effective semialgebraic) set of $kn\times kn$ matrices of the form (8). The set $\{X_{1}\cdots X_{k}\mid X_{1}\oplus\cdots\oplus X_{k}\in{\cal A}\}$ is semialgebraic (resp., effective semialgebraic).
2.4 Algebraic Irreducible sets
Let us consider the topological space obtained by equipping $\mathbb{R}^{m}$ with the Zariski topology (see [17], Ch. $1$). Given a non-empty subset $\cal Z$ of $\mathbb{R}^{m}$, $\cal Z$ is said to be irreducible if $\cal Z$ cannot be written as the union of two closed, i.e. algebraic, subsets of $\cal Z$, distinct from $\emptyset$ and $\cal Z$ itself. It can be proven that a non-empty algebraic subset $\cal Z$ of $\mathbb{R}^{m}$ is irreducible if and only if its vanishing ideal is a prime ideal (cf [17], Prop. 1.1.12). Such sets play a role in the theoretical developments presented later (see Sec. 6.2). Therefore, we recall a definition and some results. The next notation is useful.
Notation. The Zariski closure of a set $\cal A$ will be denoted by $\overline{\cal A}$.
Definition 5
Given non-empty algebraic sets ${\cal V}\subseteq\mathbb{R}^{n}$ and ${\cal W}\subseteq\mathbb{R}^{m}$, a map $f:{\cal V}\rightarrow~{}{\cal W}$ is said to be regular111 By following a well-established vocabulary in Automata Theory, the term regular will be also used in Sec. 4 to name the family of subsets of monoids defined by finite automata (or a morphism of algebraic sets), if there exist $p_{1},\dots,p_{m}\in\mathbb{R}[x_{1},.\ldots,x_{n}]$ (where the $x_{i}$ are indeterminates) such that, for all ${\bf x}=(x_{1},\ldots,x_{n})$ in $\cal V$
$f({\bf x})=(p_{1}({\bf x}),\dots,p_{m}({\bf x})).$ (9)
The following fact is well known (cf [17], Remark 1.3.2).
Lemma 1
If $f:{\cal V}\rightarrow{\cal W}$ is a regular map and $\cal V$ is irreducible, then the Zariski closure $\overline{f({\cal V})}$ of ${f({\cal V})}$ is also irreducible.
[…]
Proposition 5
The family of irreducible algebraic sets is closed with respect to the operations (7), (8) and the (finite) Cartesian product of sets.
Proof.
Let $\cal A$ be an irreducible algebraic set. By Proposition 1, with $\pi$ as in (7), $\pi({\cal A})$ is algebraic so $\overline{(\pi({\cal A}))}=\pi({\cal A})$ and the claim comes from Lemma 1 since $\pi$ is a regular map. For the operation (8) and Cartesian product, the fact that, given irreducible algebraic sets ${\cal A}_{1},\ldots,{\cal A}_{k}$, then so is the set ${\cal A}_{1}\oplus\cdots\oplus{\cal A}_{k}$ (resp., ${\cal A}_{1}\times\cdots\times{\cal A}_{k}$) comes from [17], Prop. 1.3.8-(a).∎∎
The proof of the next lemma is almost immediate.
Lemma 2
Let $H_{1},\ldots,H_{m}$ be subsets of an arbitrary monoid M, every one of which contains the identity $1$ of M. Then $(H_{1}\cup\ldots\cup H_{m})^{*}=(H_{1}\cdots H_{m})^{*}$.
Proof.
[…]
We now introduce an useful notation.
Notation. Given a subset  $\cal A$ of the group $O_{m}$ of orthogonal matrices (over $\mathbb{R}$) of order $m\geq 1$, we write ${\cal A}\in({\cal H})$ if
${\cal A}=\bigcup_{i=1}^{k}{\cal A}_{i},$ (10)
where, for every $i=1,\ldots,k$, ${\cal A}_{i}$ is an irreducible, effective algebraic set that contains the identity matrix $I$.
Proposition 6
Let $\cal A$ be a subset of $O_{m}$ with ${\cal A}\in({\cal H})$ and let $\Psi$ be a regular map
$\Psi:O_{m}\;\;\displaystyle\mathop{\longrightarrow}\;\;O_{n}.$
If, for every $i=1,\ldots,k$, $\Psi({\cal A}_{i})$ contains $I$, then $\overline{\Psi({\cal A})^{*}}\in({\cal H})$.
Proof.
Let ${\cal A}=\bigcup_{i=1}^{k}{\cal A}_{i}$ be a set as in (10). We may suppose that $k=2$, i.e. ${\cal A}={\cal A}_{1}\cup{\cal A}_{2}$, since the general case can be treated similarly. Hence $\Psi({\cal A})=\Psi({\cal A}_{1}\cup{\cal A}_{2})$ so, by Lemma 2,
$\Psi({\cal A})^{*}=(\Psi({\cal A}_{1})\Psi({\cal A}_{2}))^{*}.$
Let $H_{1}=\overline{\Psi({\cal A}_{1})\Psi({\cal A}_{2})}$. By Lemma 1, $\overline{\Psi({\cal A}_{i})}$   $(i=1,2),$ is an algebraic irreducible set. Further, by hypothesis, $\overline{\Psi({\cal A}_{i})}$ contains $I$. Following [15], Sec. 3.1, by using Gröbner bases techniques, since ${\cal A}_{i}$ is effective algebraic and $\Psi$ is a regular map, then one can effectively compute $\overline{\Psi({\cal A}_{i})}$.
Now we observe that
$\overline{\overline{{\Psi(\cal A}_{1})}\cdot\overline{\Psi({\cal A}_{2}})}\ =\ \overline{\Psi({\cal A}_{1})\cdot\Psi({\cal A}_{2})}.$ (11)
Indeed, taking the map m$:O_{n}\times O_{n}\rightarrow O_{n}$ defined, for every $X,Y\in O_{n}$, as m$(X,Y):=X\cdot Y$, one has that m is a regular and thus continuous map (w.r.t. $O_{n}$ endowed with the Zariski topology and $O_{n}\times O_{n}$ with the induced product topology). This implies, for arbitrary sets ${\cal X}_{1},{\cal X}_{2},$ that $\overline{\mbox{m}(\overline{{\cal X}_{1}},\overline{{\cal X}_{2}})}=\overline{\mbox{m}({\cal X}_{1},{\cal X}_{2})},$ and thus (11).
Let $H_{2}=\overline{H_{1}\cdot H_{1}}$. Since $I\in H_{1}$ then one has $I\in H_{2}$. By the very same argument used for $H_{1}$ one has that $H_{2}$ is an irreducible effective algebraic set such that $H_{2}=\overline{\Psi({\cal A})^{2}}$.
Finally, let us define recursively the family of sets of $O_{n}$$\{H_{i}\}_{i\geq 1},$ where $H_{0}=\{I\}$ and, for every $i\geq 1$,
$H_{i}=\overline{H_{i-1}\cdot H_{1}}.$
By proceeding as in the previous cases, one has that, for every $i\geq 3$, $H_{i}$ is an effective algebraic irreducible set such that $H_{i}=\overline{(\Psi({\cal A}_{1})\Psi({\cal A}_{2}))^{i}}$. and containing $I$. Moreover, since, for each $i\geq 1$, $I\in H_{i}$, then one has the following ascending chain
$H_{1}\;\subseteq\;H_{2}\;\subseteq\;\cdots\;H_{i}\;\subseteq\;\cdots$ (12)
which, by the irreducibility of all of its terms, terminates, i.e., there exists some $d\in\mathbb{N}$ such that
[…]
Finally let ${\cal X}=\bigcup_{i\geq 1}\ H_{i}$. By (13), $\overline{\cal X}=H_{1}\cup\cdots\cup H_{d}$. On the other hand, since, for every $i\geq 1$, $H_{i}=\overline{(\Psi({\cal A}_{1})\Psi({\cal A}_{2}))^{i}}$, then it is readily checked that
$\overline{\cal X}=H_{1}\cup\cdots\cup H_{d}=\overline{\Psi({\cal A})^{*}}.$
This proves the claim. ∎∎
Remark 2
[…]
3 Effectiveness issues
We now return to the $(L,{\cal Q})$ Intersection problem as defined in the Introduction. The main bulk of this decidability result is based upon the following proposition that is obtained by proceeding as in [11]. For the sake of completeness, we recall the proof.
Proposition 7
([10], Proposition 1) Let ${\cal Q}$ be a rational quantum automaton. Let $L\subseteq\Sigma^{*}$ be a formal language such that the set ${\mathbf{C}l}(\varphi(L))$ is effective semialgebraic. It is recursively decidable whether or not $L\cap|{\cal Q}_{>}|=\emptyset$ holds.
Proof.
Equivalently we prove the inclusion $L\subseteq|{\cal Q}_{\leq}|$, which is equivalent to
$\forall\ w\in L\Rightarrow f(w)\ =\ ||s\varphi(w)P||^{2}\ \leq\ \lambda.$
By the continuity of the map $f$, the last is equivalent to
$\forall\ X\in{\mathbf{C}l}(\varphi(L))\ \Rightarrow\ ||sXP||^{2}\ \leq\ \lambda.$ (14)
By hypothesis, there exists an effectively constructible formula $\Phi(X)$ that defines ${\mathbf{C}l}(\varphi(L))$ as a semialgebraic set (see Definition 3). Hence (14) can be written as
$\Upsilon\;\equiv\;\forall\ X:\Phi(X)\ \Rightarrow\ ||sXP||^{2}\ \leq\ \lambda.$
This implies that $L\subseteq|{\cal Q}_{\leq}|$ if and only if $\Upsilon$ holds true. Finally, one verifies whether $\Upsilon$ holds, which can be achieved by applying the Tarski Seidenberg elimination result. ∎∎
Proposition 7 shows that the decidability of the problem relies upon the computation of ${\mathbf{C}l}(\varphi(L))$. In the sequel, we will investigate the case of context-free languages. In order to deal with this case, we will use two main tools: first, an algorithm by Derksen et al. [15], already mentioned in the Introduction, that computes the Zariski closure of a finitely generated group of matrices over a field; secondly, a suitable combinatorial structuring of a context-free language based upon some special structures named cycles. The last will be presented in the next section.
4 Preliminaries on context-free languages
For the sake of self-containment and in order to fix notation, we just recall few essential terms of context-free languages and regular sets of monoids, that can be found in all introductory textbooks on theoretical computer science (see, for instance, [6, 19, 20]).
Given a monoid $M,$ a subset of $M$ is said to be regular if it is finite or it is obtained from finite sets by using finitely many times the regular operations, i.e., the union of two sets, the product ()($\cdot$)( ⋅ ) of two sets, and the star $(^{*})$, this last operation associating with a set $\cal X$ the submonoid ${\cal X}^{*}$ generated by $\cal X$. It is worth noticing that the concept of regular set can be equivalently given in terms of $M$-automaton. We will use this fact in the sequel (cf Lemma 4, Proposition 10).
A context-free grammar $G$ is a quadruple $\langle V,\Sigma,P,S\rangle$ where $\Sigma$ is the alphabet of terminal symbols, $V$ is the set of variables or nonterminal symbols, $P$ is the set of rules or productions, and $S$ is the axiom of the grammar. A word over the alphabet $\Sigma$ is called terminal. As usual, the nonterminal symbols are denoted by uppercase letters as, for instance, $A$, $B$. An arbitrary rule of the grammar is written as $A\rightarrow\alpha$, where $\alpha$ is a finite sequence of variables and terminals. The derivation relation of $G$ is denoted by $\mathop{\Rightarrow}^{*}$. The language generated by $G$ is the subset of terminal words $\{w\in\Sigma^{*}:S\displaystyle\mathop{\Rightarrow}^{*}w\}$ denoted by $L(G)$. The family of context-free languages is denoted by CFL.
Following [6], Sec. VII.5, we say that a context-free language $L$ is of index $k$, if there exist a context-free grammar $G$ with $L=L(G)$ and an integer $k\in\mathbb{N}$ such that the following property holds: for every word $w\in L(G)$, there exists some derivation $\delta=(S\displaystyle\mathop{\Rightarrow}^{*}w)$ such that the number of occurrences of variables in each sentential form of $\delta$ does not exceed $k$. The family of context-free languages of index $k$ will be denoted by $Ind(k)$ and ${\cal L}(IND_{FIN})$ will denote the family of all the languages of finite index, that is, $\bigcup_{k\in\mathbb{N}}Ind(k)$. One immediately sees that the most simple languages in such a family are those of $LIN$, the linear languages, i.e. those generated by grammars where the right hand side $\alpha$ of every rule $A\rightarrow\alpha$, contains at most one occurrence of nonterminal symbols, that is, $\alpha\in\Sigma^{*}\cup\Sigma^{*}V\Sigma^{*}$. A well-known result proven by Salomaa in [39] shows that Dyck languages cannot be generated by finite index grammars, thus implying ${\cal L}(IND_{FIN})\subset$CFL. The importance of the family ${\cal L}(IND_{FIN})$ is due to the fact that, by a well-known result by Baron and Kuich [3] (up to a technical restriction) the unambiguous context-free languages whose characteristic series in commutative variables are rational, are exactly those in ${\cal L}(IND_{FIN})$. For this reason, such languages are also called quasi-rational (see Section 4.2).
4.1 Structuring of CFL
We now find convenient to recall a combinatorial structuring of an arbitrary context-free language (see [10], Sec. 3). Such decomposition follows a recursive scheme proposed by Incitti in [28] based upon the concept defined below.
Definition 6
With each nonterminal symbol $A\in V$ associate the subset of $\Sigma^{*}\times\Sigma^{*}$ defined as
$C_{A}=\{(\alpha,\beta)\in\Sigma^{*}\times\Sigma^{*}:A\displaystyle\mathop{\Rightarrow}^{*}\alpha A\beta\}.$
$C_{A}$ will be called the set of cycles of $A$ [^2] According to the terminology of [19], a cycle is a derivation of the form $A\displaystyle\mathop{\Rightarrow}^{*}\alpha A\beta$. With a minor abuse of language, we will call cycle the pair of contexts $(\alpha,\beta)$ also
It is convenient to define the sandwich operation (cf Sec. 2.3) also for languages in the following way. If $C_{A}$ is the set above and $L^{\prime}$ an arbitrary language, we define the subset of $\Sigma^{*}$
$C_{A}\diamond L^{\prime}=\{uwv\mid(u,v)\in{C_{A}}\mbox{ and }w\in{L^{\prime}}\}.$ (15)
As the construction of the structuring mentioned above proceeds by induction on the number of nonterminal symbols, we need to show how to recombine a grammar from simpler ones obtained by choosing an arbitrary non-axiom symbol as the new axiom and by removing all the rules involving $S$. This is the reason for introducing the next notation.
Definition 7
Let $G=\langle V,\Sigma,P,S\rangle$ be a context-free grammar. Set $V^{\prime}=V\setminus\{S\}$.
For every $A\in V^{\prime}$, define the context-free grammar $G_{A}=\langle V^{\prime},\Sigma,P_{A},A\rangle$ where the set $P_{A}$ consists of all the rules $B\rightarrow\gamma$ of $G$ of the form
$\quad B\in V^{\prime},\quad\gamma\in(V^{\prime}\cup\Sigma)^{*}$
and denote by $L_{A}$ the language of all terminal words generated by the grammar $G_{A}$.
The next definition introduces the language of terminal words obtained in a derivation where $S$ occurs at the start only.
Definition 8
Let $L^{\prime}(G)$ denote the set of all the words of $\Sigma^{*}$ which admit a derivation
$S\Rightarrow\gamma_{1}\Rightarrow\cdots\Rightarrow\gamma_{\ell}\Rightarrow w$ (16)
where, for every $i=1,\ldots,\ell$, $\gamma_{i}\in(V^{\prime}\cup\Sigma)^{*}$.
The language $L^{\prime}(G)$ can be easily expressed in terms of the languages $L_{A}$ for all $A\in V^{\prime}$. Indeed, consider the set of all rules of the grammar $G$ of the form
$S\rightarrow\beta,\quad\beta\in(V^{\prime}\cup\Sigma)^{*}$ (17)
Factorize every such $\beta$ as
$\beta=w_{1}A_{1}w_{2}A_{2}\cdots w_{\ell}A_{\ell}w_{j_{\ell+1}}$ (18)
where $w_{1},\ldots,w_{\ell+1}\in\Sigma^{*}$ and $A_{1},A_{2},\ldots,A_{\ell}\in V^{\prime}$.
Proposition 8
([10], Prop. 20) The language $L$ is a finite union of languages of the form $C_{S}\diamond L^{\prime\prime}$ where
$L^{\prime\prime}=w_{1}L_{A_{1}}w_{2}L_{A_{2}}\cdots w_{\ell}L_{A_{\ell}}w_{\ell+1},$
when $\beta$ ranges over all rules (17), and with respect to the notation (18).
4.2 Structuring of ${\cal L}(IND_{FIN})$
We now recall a characterization, by Ginsburg and Spanier [18] (see also Nivat [34]), of languages of ${\cal L}(IND_{FIN})$, that will be used in Section 6.2. To this purpose, we first recall that, given alphabets $\Sigma_{1}$ and $\Sigma_{2}$, and a family $\cal F$ of languages of $\Sigma_{2}^{*}$, a $\cal F$-substitution is a morphism $\theta:\Sigma_{1}^{*}\rightarrow\Pi(\Sigma_{2}^{*})$ from the free monoid $\Sigma_{1}^{*}$ into the multiplicative monoid $\Pi(\Sigma_{2}^{*})$ of subsets of the free monoid $\Sigma_{2}^{*}$ such that
$\forall\ x\in\Sigma_{1}\quad\theta(x)\in\cal F.$
Let us define, recursively, the family $\{Qrt(k)\}_{k\in\mathbb{N}}$ of quasi-rational languages of rank $k$ (or bounded derivation languages of rank $k$) as:
$Qrt(k)=\left\{\begin{array}[]{ll}LIN&\quad\mbox{if}\ k=1,\\ LIN\circ Qrt(k-1)&\quad\mbox{if}\ k>1,\end{array}\right.$ (19)
where $LIN\circ Qrt(k-1)$ denotes the family of all the languages obtained as the images of linear languages, via $Qrt(k-1)$-substitutions, $\ k>1$, i.e.
$\{\theta(L):L\in LIN,\,\quad\forall\;x\in\Sigma_{1}\;\;\;\theta(x)\in Qrt(k-1)\}.$
The family $Qrt(k)$ is equivalently formulated in terms of the well known construct of composition of grammars (such construct is recalled in detail in the Appendix). The following holds.
Theorem 4
([18], Theorem 4.2, cf [6], Sec. VII.5, Theorem 5.2) Let $L$ be a context-free language and $k\geq 1$. Then $L\in Ind(k)$ if and only if $L\in Qrt(k)$. Moreover, $L\in Qrt(k)$ if and only if there exists a grammar $G$ with $L=L(G)$ such that $G$ is a composition ${\cal G}_{1}\circ\cdots\circ{\cal G}_{k}$ of $k$ families of linear grammars.
Example 1
Let $\cal L$ be the language over $\Sigma_{2}=\{a,b\}$ given by
=${\cal L}=\{a^{n}b^{n}:n\in\mathbb{N}\}^{*}.$
Observe that $\cal L$ is the image of the language $\{x^{n}:n\in\mathbb{N}\}$, $\Sigma_{1}=\{x\}$, under the substitution $\theta:\Sigma_{1}^{*}\rightarrow\Pi(\Sigma_{2}^{*})$, with $\theta(x)=\{a^{n}b^{n}:n\in\mathbb{N}\}$. Hence $\cal L$ is a context-free language of index $2$. Then $\cal L$ is generated by the composition $G={\cal G}_{1}\circ{\cal G}_{2}$ of the linear grammars ${\cal G}_{1}=\langle V_{1},\Sigma_{1},P_{1},S\rangle$ and ${\cal G}_{2}=\langle V_{2},\Sigma_{2},P_{2},\sigma\rangle$ where
$V_{1}=\{S\},\;\Sigma_{1}=\{\sigma\},\;P_{1}=\{S\rightarrow\varepsilon,\;S\rightarrow\sigma S\},$
and
$V_{2}=\{\sigma\},\;\Sigma_{2}=\{a,b\},\;P_{2}=\{\sigma\rightarrow\varepsilon,\;\sigma\rightarrow a\sigma b\}.$
5 Context-free languages and semialgebraic sets
Now we find convenient to recall (fixing some details of the proof) a result (cf [10], Prop. 7), obtained from the decomposition of Section 4.1. Such a result is relevant in this study since it states, for an arbitrary context-free language $L$, that the closure ${\mathbf{C}l}(\varphi(L))$ of the image of $L$, under the morphism (4) associated with a quantum automaton, is semialgebraic. To this purpose, the following construction is crucial. With an arbitrary nonterminal $A$ of $G$, associate the set $M_{A}$ defined as
$M_{A}=\{\varphi(u_{1})\oplus\varphi(u_{2})^{\scriptstyle T}\mid(u_{1},u_{2})\in C_{A}\},$ (20)
where $\varphi(u_{2})^{\scriptstyle T}$ denotes the transpose of $\varphi(u_{2})$. $M_{A}$ will be called the monoid of cycles of $A$ since the following holds.
Lemma 3
Let $A$ be a nonterminal symbol of $G$. $M_{A}$ is a monoid and its closure ${\mathbf{C}l}(M_{A})$ is an algebraic group.
Proof.
Observe first that $M_{A}$ is a subsemigroup of the group $O_{n}\oplus O_{n}.$ Indeed, if $\varphi(u_{1})\oplus\varphi(u_{2})^{\scriptstyle T}$ and $\varphi(v_{1})\oplus\varphi(v_{2})^{\scriptstyle T}$ are in $M_{A}$ then we have
$(\varphi(u_{1})\oplus\varphi(u_{2})^{\scriptstyle T})(\varphi(v_{1})\oplus\varphi(v_{2})^{\scriptstyle T})=\varphi(u_{1})\varphi(v_{1})\oplus\varphi(u_{2})^{\scriptstyle T}\varphi(v_{2})^{\scriptstyle T},$ (21)
where $(u_{1},u_{2}),(v_{1},v_{2})\in C_{A}$. On the other hand, $(u_{1}v_{1},v_{2}u_{2})\in C_{A}$ and the corresponding matrix of $M_{A}$
$\varphi(u_{1}v_{1})\oplus\varphi(v_{2}u_{2})^{\scriptstyle T}$ (22)
equals (21) for $\varphi(v_{2}u_{2})^{\scriptstyle T}=\varphi(u_{2})^{\scriptstyle T}\varphi(v_{2})^{\scriptstyle T}.$ Since $(\varepsilon,\varepsilon)\in C_{A}$, then $M_{A}$ is a monoid. The second statement comes from Theorems 2 and 3. ∎∎
Proposition 9
Let $L$ be a context-free language. The set ${\mathbf{C}l}(\varphi(L))$ is semialgebraic. Moreover, if for every $A\in V$, ${\mathbf{C}l}(M_{A})$ is effective algebraic, then ${\mathbf{C}l}(\varphi(L))$ is effective semialgebraic.
Proof.
By Proposition 8, the language $L$ is a finite union of languages of the form $C_{S}\diamond L^{\prime\prime}$ with
$L^{\prime\prime}=w_{1}L_{A_{1}}w_{2}L_{A_{2}}\cdots w_{\ell}L_{A_{\ell}}w_{\ell+1}$ (23)
where, for every $11\leq i\leq\ell+$ 1, $w_{i}\in\Sigma^{*}$ and $A_{i}\in V^{\prime}$. It suffices to show by induction on the number of nonterminal symbols that the subsets
${\mathbf{C}l}(\varphi(C_{S})\diamond\varphi(L^{\prime\prime}))$ (24)
are semialgebraic. If $\mathop{\mathrm{Card}}\nolimits(V)=1$, i.e., the set of nonterminal symbols is reduced to $S$ then $L$ is reduced to $C_{S}\diamond L^{\prime}(G)$ and $L^{\prime}(G)$ is finite. We may further assume that there is a sole terminal rule $S\rightarrow w$, so that $L^{\prime}(G)=\{w\}$. By Theorem 1 one has [^3] By Proposition 1, if ${\mathbf{C}l}(M_{S})$ is algebraic, then $\{X\oplus Y^{\scriptstyle T}:X\oplus Y\in{\mathbf{C}l}(M_{S})\}$ is so
${\mathbf{C}l}(\varphi(L))=\{X\varphi(w)Y^{\scriptstyle T}\mid X\oplus Y\oplus\{\varphi(w)\}\in{\mathbf{C}l}(M_{S}\oplus\varphi(w))\}.$ (25)
On the other hand, one has
${\mathbf{C}l}(M_{S}\oplus\varphi(w))={\mathbf{C}l}(M_{S})\oplus{\mathbf{C}l}(\varphi(w))={\mathbf{C}l}(M_{S})\oplus\varphi(w)$
which, by Lemma 3 and Proposition 3, is algebraic. By (25) and Proposition 4, we conclude that ${\mathbf{C}l}(\varphi(L))$ is effective semialgebraic. The basis of the induction is thus proved. Now assume $V$ contains more than one nonterminal symbol. By Theorem 1, Corollary 1 and Eq. (23), we get
${\mathbf{C}l}(\varphi(C_{S})\diamond\varphi(L^{\prime\prime}))=\{XZY^{\scriptstyle T}\mid X\oplus Y\oplus Z\in{\mathbf{C}l}(M_{S})\oplus{\mathbf{C}l}(\varphi(L^{\prime\prime}))\}=$
$\{XZY^{\scriptstyle T}\mid X\oplus Y\oplus Z\in{\mathbf{C}l}(M_{S})\oplus{\mathbf{C}l}(\varphi(w_{1})\varphi(L_{A_{1}})\cdots\varphi(w_{\ell})\varphi(L_{A_{\ell}})\varphi(w_{\ell+1}))\}$
By Corollary 1, we have
$\begin{array}[]{ll}&{\mathbf{C}l}(\varphi(w_{1})\varphi(L_{A_{1}})\cdots\varphi(w_{\ell})\varphi(L_{A_{\ell}})\varphi(w_{\ell+1}))\\ =&\varphi(w_{1}){\mathbf{C}l}(\varphi(L_{A_{1}}))\cdots\varphi(w_{\ell}){\mathbf{C}l}(\varphi(L_{A_{\ell}}))\varphi(w_{\ell+1}))\end{array}$
which shows, via Proposition 1 and by the induction hypothesis, that this subset is semialgebraic. Then its direct sum with ${\mathbf{C}l}(M_{S})$ is semialgebraic. We conclude by applying Proposition 4. The second part of the statement is obtained from the hypothesis by following the same proof. ∎∎
Lemma 4
Assume that $L$ is a context-free linear language and the quantum automaton is rational. Then, for every variable $A$, ${\mathbf{C}l}(M_{A})$ is effective algebraic.
Proof.
[…]
Proposition 9 together with Lemma 4 finally yield.
Corollary 2
([10], Prop. 22) If $L$ is a context-free linear language and the quantum automaton $\cal Q$ is rational, then its closure ${\mathbf{C}l}(\varphi(L))$ is an effective semialgebraic set. As a consequence, it is recursively decidable whether or not $L\cap|{\cal Q}_{>}|=\emptyset$ holds.
An intermediate class of languages in between the linear and the finite index ones, is that of metalinear (or ultralinear) languages ${\cal L}(Meta_{Lin})$. By definition, a language $L$ is in ${\cal L}(Meta_{Lin})$ if there exist a context-free grammar $G$ with $L=L(G)$ and an integer $k\in\mathbb{N}$ such that, in every derivation of $G$, the number of occurrences of variables in each sentential form does not exceed $k$. The following corollary holds.
Corollary 3
If $L\in{\cal L}(Meta_{Lin})$ and $\cal Q$ is a rational quantum automaton, then it is recursively decidable whether or not $L\cap|{\cal Q}_{>}|=\emptyset$ holds,
Proof.
It is known (see [6, 19]) that $L$ is in ${\cal L}(Meta_{Lin})$ if and only if $L$ is a finite union of finite products of context-free linear languages. Thus the claim follows from Corollary 2, the closure properties of Section 2.3 and Proposition 7. ∎∎
6 Main results
We will prove now that the problem is decidable for two families of languages: the languages generated by the class of restricted matrix context-free grammars and, subsequently, the languages generated by context-free monoidal grammars.
6.1 The case of finite index matrix languages
Now we propose a first non trivial extension of the main results of decidability provided in [10]. Such an extension is based upon the machinery of Section 5, and some combinatorial tools developed by Choffrut, Varricchio and the second author in [13] for the study of the counting functions of rational relations. To this purpose, following [14], Ch. $1$, Sec. $1.1$, we first recall that a matrix context-free grammar (which we will henceforth simply call matrix grammar) is a tuple $G=\langle V,\Sigma,M,S\rangle$, where $V$, $\Sigma$, and $M$ are respectively the finite sets of nonterminals, terminals, and matrices (of rules), and $S\in V$ is the start symbol. Each matrix $m\in M$ is a finite sequence $m=(p_{1},\ldots,p_{s})$ where each $p_{i}$, $1\leq i\leq s$, is a context-free production from $V$ to $(V\cup\Sigma)^{*}$. We denote by $P_{M}$ the set of all the context-free productions that appear in the matrices of $M$. For an arbitrary $m\in M$, we define the $1$-step derivation in $G$, $x\mathrel{\mathop{\Rightarrow}_{m}}y,$ with $x,y\in(V\cup\Sigma)^{*}$ if
$x=x_{0}\mathrel{\mathop{\Rightarrow}_{p_{1}}}x_{1}\mathrel{\mathop{\Rightarrow}_{p_{2}}}\cdots\mathrel{\mathop{\Rightarrow}_{p_{s}}}x_{s}=y,$
where $\mathrel{\mathop{\Rightarrow}_{p_{i}}}$ denotes the standard context-free derivation relation, defined by the rule $p_{i}$. This is equivalent to say that $x=x_{0},x_{s}=y$, and, for every $i=1,\ldots,s,$
$x_{i-1}=x^{\prime}_{i-1}X_{i}x^{\prime\prime}_{i-1},\quad x_{i}=x^{\prime}_{i-1}\alpha_{i}x^{\prime\prime}_{i-1},$
where $p_{i}=(X_{i}\rightarrow\alpha_{i})$, and for some $x^{\prime}_{i-1},x^{\prime\prime}_{i-1}\in(V\cup\Sigma)^{*}$. In other words, a $1$-step derivation in the matrix grammar $G$ corresponds to a $s$-step derivation in the context-free grammar $\langle V,\Sigma,P_{M},S\rangle$, i.e., using, one by one, all the context-free productions of $m$ (w.r.t. the order the $p_{i}$’s appear in $m$). Let $M^{*}$ be the set of the finite sequences of matrices and $\alpha\in M^{*}$. If $\alpha=\varepsilon$, then $x\Rightarrow_{\alpha}x,x\in V^{*}$; if one has
$\alpha=m_{1}m_{2}\cdots m_{s}\mbox{\quad and\quad}x_{0}\mathrel{\mathop{\Rightarrow}_{m_{1}}}x_{1}\mathrel{\mathop{\Rightarrow}_{m_{2}}}\cdots\ x_{s-1}\mathrel{\mathop{\Rightarrow}_{m_{s}}}x_{s},$
with $x_{i}\in(V\cup\Sigma)^{*},$ $m_{j}\in M$, $1\leq j\leq s$, $0\leq i\leq s$, then we write $x_{0}\mathrel{\mathop{\Rightarrow}_{\alpha}}x_{s}$. The language generated by $G$ is $L(G)=\{w\in\Sigma^{*}\mid S\mathrel{\mathop{\Rightarrow}_{\alpha}}w\}$.
The following definition is instrumental. It corresponds to a specific case of the simple matrix grammars introduced by Ibarra in [24] (see also [27]).
Definition 9
A matrix grammar $G=(V,\Sigma,M,S)$ is said to be restricted of index $k$, or simply restricted, if the set of nonterminals of $G$ has the form $V=\{S\}\cup V^{\prime}$, where $S\notin V^{\prime}$ and $V^{\prime}=V_{1}\cup\dots\cup V_{k},$ where $V_{i}\cap V_{j}=\emptyset,$ for all $i\neq j$, and a matrix of rules of $G$ can be only of one of the following forms
  • $(X_{1}\rightarrow u_{1}Y_{1}v_{1},\ldots,X_{k}\rightarrow u_{k}Y_{k}v_{k})$, for some $X_{i},Y_{i}\in V_{i},\ u_{i},\ v_{i}\in\Sigma^{*}$, with $1\leq i\leq k$,
  • $(X_{1}\rightarrow\varepsilon,\ldots,X_{k}\rightarrow\varepsilon)$, with $X_{i}\in V_{i},$ $1\leq i\leq k$,
  • $(S\rightarrow X_{1}\cdots X_{k})$, for some $X_{i}\in V_{i}$, with $1\leq i\leq k$.
Such grammars are very akin to simple matrix linear grammars (cf [24], pag. 363, [14], Ch. $1$, Sec. $5$).
Example 2
Let $L=\{a^{n}b^{n}c^{n}:n\in\mathbb{N}\}$ be the language over the alphabet $\Sigma=\{a,b,c\}$. Let $G$ be the restricted matrix grammar with the matrices of rules $m_{1}=(S\rightarrow ABC)$, $m_{2}=(A\rightarrow aA,B\rightarrow bB,C\rightarrow cC),$ and $m_{3}=(A\rightarrow~{}\varepsilon,B\rightarrow~{}\varepsilon,C\rightarrow\varepsilon).$ Note that the other components of $G$ can be deduced from these matrices. One easily checks that $L=L(G)$. Indeed, each word $a^{n}b^{n}c^{n},n\geq 0,$ can be generated by using first the matrix $m_{1}$, then the matrix $m_{2}$ $n$ times and finally the matrix $m_{3}$. Hence $L\subseteq L(G)$. The reverse inclusion is proved similarly.
The main result of this section is the following.
Proposition 10
Let $L$ be a language generated by a restricted grammar and let $\cal Q$ be a quantum automaton. Then one has:
  1. i.
    ${\mathbf{C}l}(\varphi(L))$ is semialgebraic.
  2. ii.
    If the quantum automaton ${\cal Q}$ is rational, then ${\mathbf{C}l}(\varphi(L))$ is effective semialgebraic and the $(L,Q)$ Intersection problem is decidable.
Proof.
Let $L$ be a language generated by a grammar as in Definition 9. Let us prove (i). Set ${\cal X}=(V^{\prime})^{k}$, and consider the sets
${\cal Y}=\{X_{1}\cdots X_{k}\in{\cal X}:\exists\ m\in M:m=(S\rightarrow X_{1}\cdots X_{k})\},$
and
${\cal Z}=\{X_{1}\cdots X_{k}\in{\cal X}:\exists\ m\in M:m=(X_{1}\rightarrow\varepsilon,\ldots,X_{k}\rightarrow\varepsilon)\}.$
Denote by $(\Sigma^{*})^{2k}$ the multiplicative monoid of $2k$-fold relations on $\Sigma^{*}$. It is easily checked that, given $w\in\Sigma^{*}$, $w\in L$ if and only if there exists some $\omega=(u_{1},v_{1},u_{2},v_{2},\ldots,u_{k},v_{k})\in(\Sigma^{*})^{2k},$ such that
$w=u_{1}v_{1}u_{2}v_{2}\;\cdots\;u_{k}v_{k},$
and $\omega$ belongs to the set $\cal L$ defined as
${\cal L}=\bigcup_{p\in{\cal Y},\;q\in\cal Z}\;{\cal L}_{pq},$ (26)
where, for $p=X_{1}\cdots X_{k},\ q=Y_{1}\cdots Y_{k}$, ${\cal L}_{pq}$ is the subset of $(\Sigma^{*})^{2k}$ defined as
${\cal L}_{pq}=\{\;(u_{1},v_{1},\ldots,u_{k},v_{k})\;:\;\exists\ \alpha\in M^{*}:\ X_{1}\cdots X_{k}\Rightarrow_{\alpha}u_{1}Y_{1}v_{1}\cdots u_{k}Y_{k}v_{k}\;\}.$
Let $s=(q_{1},\ldots,q_{n})$ be a repetition-free sequence of elements of $\cal X$, i.e., for every positive integers $i,j$, if $i\neq j$, then $q_{i}\neq q_{j}$. Clearly, $n$ is less than or equal to the cardinal number of $\cal X$. Denote by $\cal S$ the set of all such repetition-free sequences and observe that this set is finite. Further, for every $p,q\in\cal X$, let ${\cal S}_{pq}$ be the set
${\cal S}_{pq}=\{s\in{\cal S}\;\mid\;s=(q_{1},\ldots,q_{n}),\;q_{1}=p,\;q_{n}=q\}.$
Now, for every $p=X_{1}\cdots X_{k},q=Y_{1}\cdots Y_{k}$, define the subset $E_{pq}$ of $(\Sigma^{*})^{2k}$ as:
$E_{pq}=\{\;(u_{1},v_{1},\ldots,u_{k},v_{k})\;:\;\exists\ m\in M:\ X_{1}\cdots X_{k}\Rightarrow_{m}u_{1}Y_{1}v_{1}\cdots u_{k}Y_{k}v_{k}\;\},$
and the subset $C_{p}$ of $(\Sigma^{*})^{2k}$ as:
$C_{p}=\{\;(u_{1},v_{1},\ldots,u_{k},v_{k})\;:\;\exists\ \alpha\in M^{*}:\ X_{1}\cdots X_{k}\Rightarrow_{\alpha}u_{1}X_{1}v_{1}\cdots u_{k}X_{k}v_{k}\;\}.$
For every $s=(q_{1},\ldots,q_{n})\in\cal S$, define the (possibly empty) subset ${\cal L}_{s}$ of $(\Sigma^{*})^{2k}$ as:
${\cal L}_{s}=C_{q_{1}}\diamond E_{q_{1}q_{2}}\diamond C_{q_{2}}\diamond\cdots\diamond C_{q_{n-1}}\diamond E_{q_{n-1}q_{n}}\diamond C_{q_{n}},$ (27)
where ()($\diamond$)( ⋄ ) corresponds to the extension of the operation (15) to $(\Sigma^{*})^{2k}$, that is, for subsets ${\cal M},{\cal M}^{\prime}$ of $(\Sigma^{*})^{2k}$, the tuple $(w_{1},\ldots,w_{2k})\in{\cal M}\diamond{\cal M}^{\prime}$ if and only if there exist
$(u_{1},v_{1},\ldots,u_{k},v_{k})\in{\cal M}\;\;\mbox{ and }\;\;(u^{\prime}_{1},v^{\prime}_{1},\ldots,u^{\prime}_{k},v^{\prime}_{k})\in{\cal M}^{\prime}$
such that, for every $i=1,\ldots,k,$
$w_{2i-1}\ =\ u_{i}u^{\prime}_{i}\;\;\;\;\mbox{ and }\;\;\;\;w_{2i}=v^{\prime}_{i}v_{i}.$
By using an argument similar to that of [13], Theorem 2, one checks that
${\cal L}_{pq}=\bigcup_{s\in{\cal S}_{pq}}\ {\cal L}_{s}.$ (28)
Let now consider the monoid morphism
$\widehat{\varphi}:(\Sigma^{*})^{2k}\rightarrow O_{2nk},$
which maps every element $\omega=(u_{1},v_{1},\ldots,u_{k},v_{k})$ into the orthogonal matrix
$\widehat{\varphi}(\omega)=\varphi(u_{1})\ \oplus\ \varphi(v_{1})\ \oplus\ \cdots\ \oplus\ \varphi(u_{k})\ \oplus\ \varphi(v_{k}).$
Since, by Theorem 1, one has
${\mathbf{C}l}({\varphi}(L))\ =\ \{M_{1}\cdots M_{2k}\ :M_{1}\oplus\cdots\oplus M_{2k}\in{\mathbf{C}l}(\widehat{\varphi}(\cal L))\},$
to show that ${\mathbf{C}l}({\varphi}(L))$ is semialgebraic, by Proposition 4, it is enough to show that ${\mathbf{C}l}(\widehat{\varphi}(\cal L))$ is so. To this purpose, with an arbitrary $q\in{\cal X}$, associate the subset $M_{q}$ of $O_{2nk}$ defined as
$\{\varphi(u_{1})\oplus\varphi(v_{1})^{\scriptstyle T}\oplus\cdots\oplus\varphi(u_{k})\ \oplus\ \varphi(v_{k})^{\scriptstyle T}\mid(u_{1},v_{1},\ldots,u_{k},v_{k})\in C_{q}\}.$ (29)
By proceeding as in Lemma 3, one proves that $M_{q}$ is a monoid and its closure ${\mathbf{C}l}(M_{q})$ is an algebraic group. From this, by Proposition 1, it follows that ${\mathbf{C}l}(\widehat{\varphi}({C}_{q}))$ is an algebraic set. This implies, via Theorem 1 and Proposition 2, that every set ${\mathbf{C}l}(\widehat{\varphi}({\cal L}_{s}))$ of (27) is semialgebraic. Then the semialgebraicity of ${\mathbf{C}l}(\widehat{\varphi}(\cal L))$ follows, by applying Proposition 1, from the last condition and (26), (28).
(ii) Assume that the quantum automaton ${\cal Q}$ is rational. First we prove that ${\mathbf{C}l}(M_{q})$ is effective algebraic. One proceeds as in the proof of Lemma 4. Indeed, since $L$ is generated by a restricted matrix grammar, then $M_{q}$ is a regular submonoid of $O_{2nk}$. Precisely, it is recognized by the finite $O_{2nk}$–automaton whose states are the elements of $\cal X$, the initial and final states coincide with $q$, and the transitions are of the form
[…]
where, being $s=X_{1}\cdots X_{k},\ t=Y_{1}\cdots Y_{k}$, there exists $m\in M$ such that
$X_{1}\cdots X_{k}\Rightarrow_{m}u_{1}Y_{1}v_{1}\cdots u_{k}Y_{k}v_{k},$
with $\ell(e)\in O_{2nk}$ defined as
$\ell(e)=\varphi(u_{1})\oplus\varphi(v_{1})^{\scriptstyle T}\oplus\cdots\oplus\varphi(u_{k})\ \oplus\ \varphi(v_{k})^{\scriptstyle T}.$
Hence $M_{q}$ is a regular monoid and the group $\langle M_{q}\rangle$ has an effective finite generating set. By applying to $\langle M_{q}\rangle$ the algorithm of Derksen et al. one gets an effective presentation of ${\mathbf{C}l}(M_{q})$. From this, by Proposition 1, it follows that ${\mathbf{C}l}(\widehat{\varphi}({C}_{q}))$ is effective algebraic as well. Finally, since all the steps of the proof (i) are effective, the fact that ${\mathbf{C}l}(\widehat{\varphi}({C}_{q}))$ is effective algebraic implies that ${\mathbf{C}l}({\varphi}(L))$ is effective semialgebraic as well. The claim follows by applying Proposition 7. ∎∎
We now discuss some straightforward corollaries of Proposition 10. First, remark that, by applying the proposition to the matrix grammars of index $k=1$, we recover [10], Prop. 22 (cf Corollary 2). Moreover, note that the language of Example 2 is an instance of the following more general result. To this purpose, we recall that a bounded semilinear language is a finite union of linear languages which are languages of the form
$L=\{u_{1}^{n_{1}}\cdots u_{k}^{n_{k}}\mid(n_{1},\ldots,n_{k})\in{R}\},$ (30)
for some fixed words $u_{i}\in\Sigma^{*}$ for $i=1,\ldots,k$ and $R\subseteq\mathbb{N}^{k}$ is a linear set, i.e., there exists ${\bf v}_{0},{\bf v}_{1},\ldots,{\bf v}_{\ell}\in\mathbb{N}^{k}$ such that
$R=\{{\bf v}_{0}+\lambda_{1}{\bf v}_{1}+\cdots+\lambda_{\ell}{\bf v}_{\ell}\mid\lambda_{1},\ldots,\lambda_{\ell}\in\mathbb{N}\}.$ (31)
Lemma 5
Every bounded semi-linear language can be generated by a restricted grammar.
Proof.
It is enough to prove the claim for an arbitrary language $L$ of the form (30). It is checked that $L$ is generated by the restricted grammar of index $k$ $G=\langle V,\Sigma,M,S\rangle$, where $V=\{S,X_{1},\ldots,X_{k}\}$ and $M$ is given by the matrices
  • $(S\rightarrow X_{1}\cdots X_{k})$,
  • $(X_{1}\rightarrow\varepsilon,\ldots,X_{k}\rightarrow\varepsilon)$,
  • $(S\rightarrow u_{1}^{{\bf v}_{01}}X_{1}\cdots u_{k}^{{\bf v}_{0k}}X_{k})$, with ${{\bf v}_{0}=({\bf v}_{01}},\ldots,{\bf v}_{0k})$,
  • For every $i=1,\ldots,\ell,$ $(X_{1}\rightarrow u_{1}^{{\bf v}_{i1}}X_{1},\ldots,X_{k}\rightarrow u_{k}^{{\bf v}_{ik}}X_{k})$
[…]
where the vectors ${\bf v}_{i},\ i=0,\ldots,k$, define (31). ∎
Proposition 10 and Lemma 5 provide a new proof of the following.
Corollary 4
([10], Prop. 21) If $L$ is bounded semilinear then its closure ${\mathbf{C}l}(\varphi(L))$ is semialgebraic. Furthermore, if the quantum automaton ${\cal Q}$ is rational, then the $(L,Q)$ Intersection problem is decidable.
Example 2 (continued). Let ${\cal Q}=(s,\varphi,P,\lambda)$ be a finite quantum automaton where $\varphi:\Sigma^{*}\longrightarrow O_{n}$ is the morphism (4) associated with ${\cal Q}$. Recall that $\mathbf{C}l{({\cal A})}$ and $\overline{{\cal A}}$ denote, respectively, the Euclidean and the Zariski closure of a set $\cal A$. The set ${\mathbf{C}l}(\varphi(L))$ is then equal to
${\mathbf{C}l}(\{\varphi(a)^{n}\varphi(b)^{n}\varphi(c)^{n}:n\in\mathbb{N}\})={\mathbf{C}l}(\Psi(\{\varphi(a)^{n}\oplus\varphi(b)^{n}\oplus\varphi(c)^{n}:n\in\mathbb{N}\}))$ (32)
where ()direct-sum($\oplus$)( ⊕ ) is the operation defined in (8) and $\Psi$ is the map defined, for every $X,Y,Z\in O_{n}$, as $\Psi(X\oplus Y\oplus Z):=XYZ$. By Theorem 1, the right-side term of (32) rewrites as $\Psi({\mathbf{C}l}(\{\varphi(a)^{n}\oplus\varphi(b)^{n}\oplus\varphi(c)^{n}:n\in\mathbb{N}\}))$, so that
${\mathbf{C}l}(\varphi(L))=\Psi({\mathbf{C}l}(\{g\}^{*}))=\Psi({\mathbf{C}l}(\langle\{g\}\rangle)),$
where $g=\varphi(a)\oplus\varphi(b)\oplus\varphi(c)$. By Theorem 3, ${\mathbf{C}l}(\langle\{g\}\rangle)=\overline{{\langle\{g\}\rangle}}$ is algebraic. From this, by Proposition 4, it follows that ${\mathbf{C}l}(\varphi(L))=\Psi(\overline{{\langle\{g\}\rangle}})$ is semialgebraic.
If the automaton $\cal Q$ is a rational, then, by applying Derksen’s algorithm to $\{g\}$, one computes $\overline{{\langle\{g\}\rangle}}$ as an algebraic set, which, in turn, implies that $\Psi(\overline{{\langle\{g\}\rangle}})$ is effective semialgebraic.
We observe that, not only, Proposition 10 provides a unitary scheme for the decidability results of [10] (i.e., Corollary 2 and Corollary 4) but it gives a non trivial extension of them. Indeed the following example provides languages, neither bounded nor context-free, defined by restricted grammars, for which the Intersection Problem is now decidable.
Example 3
Let $L_{k}$, with $k\in\mathbb{N}$, be the language over $\Sigma=\{a,b\}$ given by $L_{k}=\{u^{k}:u\in\Sigma^{*}\}$. It is easily checked that $L_{k}$ can be generated by the restricted grammar $G=\langle V,\Sigma,M,S\rangle$, where $V=\{S,X_{1},\ldots,X_{k}\}$, and $M$ is given by the matrix rules $(S\rightarrow X_{1}\cdots X_{k})$, $(X_{1}\rightarrow\varepsilon,\ldots,X_{k}\rightarrow~{}\varepsilon)$, and, for every $\sigma\in\Sigma$, $(X_{1}\rightarrow\sigma X_{1},\ldots,X_{k}\rightarrow\sigma X_{k})$. Note that the very same argument used in Example 2 allows one to verify that ${\mathbf{C}l}(\varphi(L))$ is semialgebraic and effective semialgebraic in the case that the quantum automaton is rational.
6.2 The case of monoidal languages
We start with the following definition that is instrumental for this section. We first recall that a context-free linear grammar is said to be minimal linear if it has a unique variable.
Definition 10
A grammar $G=\langle V,\Sigma,P,S\rangle$ is said to be monoidal of index $k$ if it is the composition
$G={\cal G}_{1}\circ{\cal G}_{2}\circ\cdots\circ{\cal G}_{k},$
of $k$ families of minimal linear grammars where every terminal production $\ (X\rightarrow u),\ u\in\Sigma^{*},$ is such that $u=\varepsilon$.
A language is said to be monoidal if it is generated by a monoidal grammar of index $k$, for some $k\in\mathbb{N}$. The main result of this section is the following.
Proposition 11
Let $(L,Q)$ be a pair where $Q$ is a rational quantum automaton and $L$ is a language generated by a monoidal grammar $G={\cal G}_{1}\circ{\cal G}_{2}\circ\cdots\circ{\cal G}_{k}$. If the Zariski closures of the monoids of cycles of the grammars of ${\cal G}_{k}$ – i.e., the lowest level of $G$ – are irreducible, then ${\mathbf{C}l}(\varphi(L))$ is effective semialgebraic.
Remark 3
Examples of monoids generated by a finite number of matrices whose Zariski closure is irreducible can be borrowed from [16]. In Theorem 1.1 of loc.cit., the authors determine the Zariski closure of a cyclic matrix semigroup over $\mathbb{C}$. By the proof [22, Corollary 17], the above result implies that if we consider the cyclic matrix semigroup generated by a real matrix with eigenvalues greater than $1$, its Zariski closure is irreducible over $\mathbb{R}$. Similarly, Example 3.8 from [16] can be adapted to produce an example of a matrix semigroup $M$, generated by two matrices, such that the Zariski closure of $M$ is irreducible over $\mathbb{R}$.
We will do the proof of Proposition 11 for $k\leq 2$, i.e., for a monoidal grammar $G={\cal G}_{1}\circ{\cal G}_{2}$ of index $2$, since, by (19), the general case is treated with the same argument. The proof is technically structured in the following lemmata. Also we find useful to recall a notation introduced in Section 2.4.
Notation. Given a subset $\cal A$ of the group $O_{m}$ of orthogonal matrices (over $\mathbb{R}$), we write
${\cal A}\in({\cal H})$ (33)
if $\cal A$ is a finite union of irreducible, effective algebraic sets, every one of which contains the identity matrix $I$ of $O_{m}$.
Lemma 6
Let $G={\cal G}_{1}$ be a monoidal grammar of index $1$ and let $L=L(G)$. Under the assumption of Proposition 11, there exist a set $H\in({\cal H})$ and a regular map $\Psi$ such that ${\mathbf{C}l}(\varphi(L))=\Psi(H).$ Moreover ${\mathbf{C}l}(\varphi(L))$ is effective semialgebraic and contains $I$.
Proof.
Set $G=\langle V,\Sigma,P,S\rangle$ and observe that $V=\{S\}$, so that $G$ has a sole monoid of cycles $M_{S}$. Taking into account (25) and the form of the terminal productions of $G$, one has ${\mathbf{C}l}(\varphi(L_{S}))=\Psi({\mathbf{C}l}(M_{S}))$, where $\Psi(X\oplus Y):=XY^{\scriptstyle T}.$ From this and the fact that, by hypothesis, ${\mathbf{C}l}(M_{S})$ is irreducible, the claim follows. ∎∎
Now assume $G={\cal G}_{1}\circ{\cal G}_{2}$ and let ${\cal G}_{1}=\langle V_{1},\Sigma_{1},P_{1},S\rangle$ be the first grammar of the composition $G$. Set $V_{1}=\{S\}$. In the sequel, for every $A\in\Sigma_{1}$, we denote by $L_{A}$ the language generated by the grammar $G_{A}\in{\cal G}_{2}$. Let $\Sigma_{1}^{*}\times\Sigma_{1}^{*}$ be the multiplicative monoid of binary relations on $\Sigma_{1}^{*}$ and let $\Pi(O_{n}\oplus O_{n})$ be the multiplicative monoid of subsets of $O_{n}\oplus O_{n}$. Define now the morphism between the two monoids
$\zeta:\ \Sigma_{1}^{*}\times\Sigma_{1}^{*}\ \longrightarrow\ \Pi(O_{n}\oplus O_{n})$
as: for every $(u,v)\in\Sigma_{1}^{*}\times\Sigma_{1}^{*}$, with $u=A_{1}\cdots A_{s},\ v=B_{1}\cdots B_{t},\ A_{i},B_{i}\in\Sigma_{1}$
$\zeta(u,v)\ =\ (\varphi(L_{A_{1}})\cdots\varphi(L_{A_{s}})\ \oplus\ \varphi(L_{B_{1}})^{\scriptstyle T}\cdots\varphi(L_{B_{t}})^{\scriptstyle T})$ (34)
where, for every $i=1,\ldots,t$, $\varphi(L_{B_{i}})^{\scriptstyle T}$ denotes the set of matrices
$\{m^{\scriptstyle T}\ :\ m\in\varphi(L_{B_{i}})\}.$ (35)
The following lemma holds.
Lemma 7
There exists an effective computable finite subset $\cal R$ of $\ \Sigma_{1}^{*}\times\Sigma_{1}^{*}$ such that $M_{S}=\zeta({\cal R}^{*})$.
Proof.
Let us first define the set
$\widehat{C}_{S}=\{(\alpha,\beta)\in\Sigma_{1}^{*}\times\Sigma_{1}^{*}:(\alpha,\beta^{\sim})\in C_{S}\},$
where $\beta^{\sim}$ denotes the mirror image of $\beta$. Observe that $\widehat{C}_{S}$ is a submonoid of $\Sigma_{1}^{*}\times\Sigma_{1}^{*}$. Moreover one has
$M_{S}=\zeta(\widehat{C}_{S}).$
Indeed, let us check $M_{S}\subseteq\zeta(\widehat{C}_{S}).$ If $\varphi(u)\oplus\varphi(v)^{\scriptstyle T}\in M_{S}$, then $S\displaystyle\mathop{\Rightarrow}^{*}uSv$. Since $G={\cal G}_{1}\circ{\cal G}_{2}$, the last derivation can be rewritten as
$S\displaystyle\mathop{\Rightarrow}^{*}\alpha S\beta$ (36)
where (36) is a derivation of ${\cal G}_{1}$ and
$\alpha\displaystyle\mathop{\Rightarrow}^{*}u,\quad\quad\beta\displaystyle\mathop{\Rightarrow}^{*}v,$ (37)
where both the derivations of (37) are in the grammars of ${\cal G}_{2}.$ Now set $\alpha=A_{1}A_{2}\cdots A_{s},$ and $\beta=B_{s}B_{s-1}\cdots B_{1}$, with $A_{i},B_{i}\in\Sigma_{1}\cup\{\varepsilon\},s\geq 0$. From (37), one then has
$u=u_{1}\cdots u_{s},\quad\quad v=v_{s}\cdots v_{1},$
where, for every $i=1,\ldots,s,$ $u_{i}\in L_{A_{i}},$ and $v_{i}\in L_{B_{i}}$. From this, one has
$\varphi(u)\in\varphi(L_{A_{1}})\cdots\varphi(L_{A_{s}}),\quad\quad\varphi(v)^{\scriptstyle T}\in\varphi(L_{B_{1}})^{\scriptstyle T}\cdots\varphi(L_{B_{t}})^{\scriptstyle T}.$ (38)
Now observe that, by (36), $(\alpha,\beta)\in C_{S}$ so that $(\alpha,\beta^{\sim})\in\widehat{C}_{S}$. By (38) the latter implies $\varphi(u)\oplus\varphi(v)^{\scriptstyle T}\in\zeta(\widehat{C}_{S})$. Hence $M_{S}\subseteq\zeta(\widehat{C}_{S})$. The other inclusion is checked similarly. Finally observe that a finite set $\cal R$ of binary relations on $\Sigma_{1}^{*}$ generating $\widehat{C}_{S}$ as a monoid, can be effectively computed from the productions of ${\cal G}_{1}$. The lemma is proved. ∎∎
Lemma 8
${\mathbf{C}l}(M_{S})\in({\cal H}).$ In particular, ${\mathbf{C}l}(M_{S})$ is effective algebraic.
Proof.
By Lemma 7, there exists an effective computable finite subset $\cal R$ of $\ \Sigma_{1}^{*}\times\Sigma_{1}^{*}$ such that $M_{S}=\zeta({\cal R}^{*})$. First let us prove the following claim:
Claim. Let $\cal R$ =$=\{r\}$, with $r\in\Sigma_{1}^{*}\times\Sigma_{1}^{*}$. Then ${\mathbf{C}l}(\zeta({\cal R}))$ contains $I$ and ${\mathbf{C}l}(\zeta({\cal R}))=\Psi(H)$ for some $H\in({\cal H})$ (cf (33)) and regular map $\Psi$.
Let us write $r=(\alpha,\beta)$ with $\alpha=A_{1}A_{2}\cdots A_{s},$ $\beta=B_{1}\cdots B_{t}$, $A_{i},B_{i}\in\Sigma_{1},\ s,t\geq 0$. We assume first that $\alpha$ and $\beta$ have the same length, i.e., $s=t$. By (34), the fact that $\zeta$ is a monoid morphism and Corollary 1, one gets
${\mathbf{C}l}(\zeta({\cal R}))=\ ({\mathbf{C}l}(\varphi(L_{A_{1}}))\cdots{\mathbf{C}l}(\varphi(L_{A_{s}}))\ \oplus\ {\mathbf{C}l}(\varphi(L_{B_{1}}))^{\scriptstyle T}\cdots{\mathbf{C}l}(\varphi(L_{B_{s}}))^{\scriptstyle T}),$
taking into account that, for an arbitrary set $\cal A$ of orthogonal matrices, for the set (35), one has ${\mathbf{C}l}({\cal A}^{\scriptstyle T})={\mathbf{C}l}({\cal A})^{\scriptstyle T}$. By Lemma 6, the previous formula gives
${\mathbf{C}l}(\zeta({\cal R}))=\ (\Psi(H_{1})\cdots\Psi(H_{s})\ \oplus\ \Psi(G_{1})^{\scriptstyle T}\cdots\Psi(G_{s})^{\scriptstyle T}),$
where $\Psi$ is a regular map and, for every $i=1,\ldots,s$,  $H_{i},G_{i}\in({\cal H})$.
Since each $\Psi(H_{i})$ and $\Psi(G_{i})$ of the last formula, contain $I$ so holds for ${\mathbf{C}l}(\zeta({\cal R}))$. Set $H=H_{1}\oplus\cdots\oplus H_{s}\oplus G_{1}\oplus\cdots\oplus G_{s}$ and observe that, by Proposition 5, $H\in({\cal H})$. Then ${\mathbf{C}l}(\zeta({\cal R}))=\ \Psi^{\prime}(H)$, where $\Psi^{\prime}$ is the regular map defined as
$\Psi^{\prime}(X_{1}\oplus\cdots\oplus X_{s}\oplus Y_{1}\oplus\cdots\oplus Y_{s})=(\Psi(X_{1})\cdots\Psi(X_{s})\oplus\Psi(Y_{1})^{\scriptstyle T}\cdots\Psi(Y_{s})^{\scriptstyle T}).$
[…]
Since $\zeta$ is a monoid morphism, then one has
${\mathbf{C}l}(M_{S})={\mathbf{C}l}(\zeta({\cal R}^{*}))={\mathbf{C}l}(\zeta({\cal R})^{*}).$
Now taking into account that, by Corollary 1, for an arbitrary set $\cal X$,
${\mathbf{C}l}({\cal X}^{*})={\mathbf{C}l}({\mathbf{C}l}({\cal X})^{*}),$ (39)
one has ${\mathbf{C}l}(M_{S})={\mathbf{C}l}({\mathbf{C}l}(\zeta({\cal R}))^{*}).$ By Theorems 2 and 3, one has
${\mathbf{C}l}({\mathbf{C}l}(\zeta({\cal R}))^{*})=\overline{{\mathbf{C}l}(\zeta({\cal R}))^{*}},$ (40)
where $\overline{{\mathbf{C}l}(\zeta({\cal R}))^{*}}$ is the Zariski closure of $({\mathbf{C}l}(\zeta({\cal R})))^{*}$. Set =${\cal R}=\{r_{1},\ldots,r_{s}\},$ with $r_{i}\in\Sigma_{1}^{*}\times\Sigma_{1}^{*},s\geq 1$. Since, for every $i=1,\ldots,s$, ${\mathbf{C}l}(\zeta({r_{i}}))$ contains the identity, by Lemma 2, one gets ${\mathbf{C}l}(\zeta({\cal R}))^{*}=({\mathbf{C}l}(\zeta({r_{1}}))\cdots{\mathbf{C}l}(\zeta({r_{s}})))^{*}$, which, thus, implies
${\mathbf{C}l}(M_{S})=\overline{({\mathbf{C}l}(\zeta({r_{1}}))\cdots{\mathbf{C}l}(\zeta({r_{s}})))^{*}}.$
By the claim, for every $i=1,\ldots,s$, ${\mathbf{C}l}(\zeta({r_{i}}))=\Psi(H_{i})$, where $H_{i}\in({\cal H})$, and $\Psi$ is a regular map not depending upon $H_{i}$. Hence one gets
${\mathbf{C}l}(M_{S})\ =\ \overline{(\Psi(H_{1})\ \cdots\ \Psi(H_{k}))^{*}}.$ (41)
Observe that in (41) we can write $\Psi(H_{1})\ \cdots\ \Psi(H_{k})$ as $\widehat{\Psi}(H_{1}\oplus\cdots\oplus H_{k})$, where $\widehat{\Psi}$ is the regular map defined as $\widehat{\Psi}(X_{1}\oplus\cdots\oplus X_{k})=\Psi(X_{1})\cdots\Psi(X_{k})$. Hence we get
${\mathbf{C}l}(M_{S})=\overline{(\widehat{\Psi}(H))^{*}},$
where, again, $H=(H_{1}\oplus\cdots\oplus H_{k})\in({\cal H})$ as one can easily check. Finally ${\mathbf{C}l}(M_{S})\in({\cal H})$ follows by applying Proposition 6. ∎∎
We are now able to prove the main result of this section.
Proof of Proposition 11: As remarked before, it is enough to prove the claim for $k\leq 2$ since, by (19), the general case is treated with the same argument. Set $G=\langle V,\Sigma,P,S\rangle$ and observe that, for every $A\in V$, the fact that ${\mathbf{C}l}(M_{A})$ is effective algebraic comes from Lemmata 4 and 8. The claim now follows from Proposition 9. ∎
By Propositions 1 and 7, immediate corollaries of Proposition 11 are the following.
Corollary 5
If the pair $(L,Q)$ satisfies the hypothesis of Proposition 11, then the $(L,Q)$ Intersection problem is decidable.
Corollary 6
Let $(L,Q)$ be a pair where $Q$ is a rational quantum automaton and $L$ is a finite union $L=\bigcup_{i\in{\cal I}}L_{i},$ of monoidal languages, every one of which is such that the corresponding pair $(L_{i},Q)$ satisfies the hypothesis of Proposition 11. Then the $(L,Q)$ Intersection problem is decidable.
7 Concluding remarks and open problems
We finally address some questions arising from this study.
[$\alpha$]  One could extend Proposition 11 to a broader class of quantum automata. In this context, the construction of the decidability procedure would require the computation of the Zariski closure of a group $G$ of matrices that is algebraically presented, i.e. $G$ admits an infinite set of generators that is algebraic itself. It may be of interest to investigate conditions ensuring the extension of the algorithms of [15, 23, 35] for such groups.
[$\beta$]  It is open whether the $(L,Q)$ Intersection problem is decidable for context-free languages. We conjecture that the answer to this problem is negative. An approach may be to reduce the problem to a so called “birdie problem” [40] (cf also [20], Ch. 8). In such a problem, assuming the feasibility of the computation of the Zariski closure of the groups above, one gets, as a consequence, the decidability of some unsolvable problem.
[$\gamma$]  The $(L,Q)$ Intersection problem is decidable for bounded semi-linear languages. It may be of interest to get extensions of this result for the broader class of languages accepted by the reversal bounded non deterministic counter machines [25]. This extension could exploit a characterization of such languages provided by Ibarra in [26] in terms of some special combinatorial systems called RLGMC-grammars (Right-Linear Grammar with Multi-Counters). A RLGMC-grammar is a regular grammar that associates with each derivation a weight, i.e. a vector in the additive monoid $\mathbb{N}^{d}$. Then a derivation is considered successful if the corresponding weight belongs to the monoid generated by $\bf 1^{d}$ $=(1,1,\ldots,1)$. In this case, the construction of the formula that makes the problem decidable, is reduced to the computation of the Zariski closure of the intersection of two finitely generated monoids of matrices, say $M_{1}$ and $M_{2}$. One can compute $\overline{M_{1}}$, $\overline{M_{2}}$ and, consequently, $\overline{M_{1}}\cap\overline{M_{2}}$. However, in general, the inclusion $\overline{M_{1}\cap M_{2}}\subseteq\overline{M_{1}}\cap\overline{M_{2}}$ is strict and the intrinsic difficulty to control the accumulation points of the set – i.e., to exclude those not coming from the effective functioning of the machines – makes such a computation a non trivial task.
Acknowledgments The authors would like to thank Mima Stanojkovski for useful discussion on her paper [16].
Appendix
We recall the construction of the context-free grammar that generates the image $\theta(L)$ of a context-free language $L$ over the alphabet $\Sigma_{1}$ under a context-free substitution $\theta:\Sigma_{1}^{*}\rightarrow\Pi(\Sigma_{2}^{*})$, i.e. a $\cal F$-substitution, with $\cal F$=CFL.
We follow verbatim the argument of [6], Sec. II.2. Let ${\cal G}_{1}=\langle V_{1},\Sigma_{1},P_{1},S\rangle$ be the context-free grammar that generates $L$ and let
${\cal G}_{2}=\{G_{x}\ :\ x\in\Sigma_{1}\}$
be the family of context-free grammars where, for every $x\in\Sigma_{1}$,
$G_{x}=\langle V_{x},\Sigma_{2},P_{x},S_{x}\rangle$
is the grammar that generates $\theta(x)$. One may assume that, for all $x\in\Sigma_{1}$, $V_{1}\cap V_{x}=\emptyset$ and, for every $x,y\in\Sigma_{1},\ x\neq y$, $V_{x}\cap V_{y}=\emptyset$. Let be the set
${\cal V}_{2}=\{S_{x}:x\in\Sigma_{1}\}$
and let the copy morphism
$\mbox{c}:(V_{1}\cup\Sigma_{1})^{*}\rightarrow(V_{1}\cup{\cal V}_{2})^{*},$ (42)
defined as: $\mbox{c}(z)=z$, if $z\in V_{1}$, and $\mbox{c}(z)=S_{z}$, if $z\in\Sigma_{1}$. With respect to c, consider the grammar $G_{\mbox{c}}$ that generates $\mbox{c}(L)$. Such a grammar is defined as
$G_{\mbox{c}}=\langle V_{1},\{S_{x}:x\in\Sigma_{1}\},P_{\mbox{c}},S\rangle,$
where $P_{\mbox{c}}=\{A\rightarrow\mbox{c}(\alpha):(A\rightarrow\alpha)\in P_{1}\}$. Finally, define
$H=\langle W,\Sigma_{2},Q,S\rangle$ (43)
as the context-free grammar where
$W=V_{1}\ \cup\ \bigcup_{x\in\Sigma_{1}}\ V_{x},\quad\quad Q=P_{\mbox{c}}\ \cup\ \bigcup_{x\in\Sigma_{1}}\ P_{x}.$
Such grammar $H$ is called the composition of ${\cal G}_{1}$ and ${\cal G}_{2}$ and denoted by
$G={\cal G}_{1}\circ{\cal G}_{2}.$ (44)
The following lemma holds.
Lemma 9
If $G$ is the grammar (44), then $\theta(L)=L({G})$.
A composition of $k$ families of context-free grammars $({\cal G}_{1},\ldots,{\cal G}_{k}),$ with $k\geq 1$ will be the context-free grammar
$G={\cal G}_{1}\circ{\cal G}_{2}\circ\cdots\circ{\cal G}_{k},$
obtained by iterating $(k-1)$ times the operation (44) of composition of grammars on the families ${\cal G}_{1},\ldots,{\cal G}_{k}$.